### 08 Dec 2023 v1

# Sharp bounds on the height of K-semistable toric Fano varieties II, the log case

Abstract.
In our previous work we conjectured - inspired by an algebro-geometric result of Fujita - that the height of an arithmetic Fano variety $\mathcal{X}$ of relative dimension $n$ is maximal when $\mathcal{X}$ is the projective space $\mathbb{P}_{\mathbb{Z}}^{n}$ over the integers, endowed with the Fubini-Study metric, if the the corresponding complex Fano variety is K-semistable. In this work the conjecture is settled for diagonal hypersurfaces in $\mathbb{P}_{\mathbb{Z}}^{n+1}.$ The proof is based on a logarithmic extension of our previous conjecture, of independent interest, which is established for toric log Fano varieties of relative dimension at most three, hyperplane arrangements on $\mathbb{P}_{\mathbb{Z}}^{n},$ as well as for general logarithmic arithmetic Fano surfaces.
1. Introduction
This is a sequel to [2], where a conjectural arithmetic analog of Fujita’s sharp bound for the degree (volume) of K-semistable Fano varieties over $\mathbb{C}$ was proposed, concerning arithmetic Fano varieties $\mathcal{X}.$ The case when $\mathcal{X}$ is the canonical integral model of a toric Fano variety $X$ was settled in [2], when the relative dimension $n$ is at most six (the extension to any $n$ is conditioned on a conjectural gap hypothesis for the algebro-geometric degree). Here we will, in particular, show that the conjecture introduced in [2] holds for any diagonal Fano hypersurface $\mathcal{X}$ in $\mathbb{P}_{\mathbb{Z}}^{n+1}$ (see Section 1.1.3 below). The proof is based on the following extension of the conjecture in [2] to the logarithmic setting, which is the main focus of the present work:
Conjecture 1.1.
Let $(\mathcal{X},\mathcal{D})$ be an arithmetic log Fano variety. Then the following inequality of arithmetic intersection numbers holds for any volume-normalized continuous metric on $-(K_{X}+\Delta)$ with positive curvature current if $(X,\Delta)$ is K-semistable:
$(\overline{-\mathcal{K}_{(\mathcal{X},\mathcal{D})}})^{n+1}\leq(\overline{-\mathcal{K}_{\mathbb{P}_{\mathbb{Z}}^{n}}})^{n+1},$
where $-K_{\mathbb{P}_{\mathbb{C}}^{n}}$ is endowed with the volume-normalized Fubini-Study metric. Moreover, if $\mathcal{X}$ is normal equality holds if and only if $(\mathcal{X},\mathcal{D})=(\mathbb{P}_{\mathbb{Z}}^{n},0)$ and the metric is Kähler-Einstein, i.e. coincides with the Fubini-Study metric up to the action of an automorphism of $\mathbb{P}_{\mathbb{C}}^{n}.$
By definition, an arithmetic log Fano variety $(\mathcal{X},\mathcal{D})$ is a projective flat scheme $\mathcal{X}$ over $\mathbb{Z}$ together with an effective $\mathbb{Q}-$divisor $\mathcal{D}$ on $\mathcal{X}$ such that
$-\mathcal{K}_{(\mathcal{X},\mathcal{D})}:=-(\mathcal{K}_{\mathcal{X}}+\mathcal{D})$
defines a relatively ample $\mathbb{Q}-$line bundle, where $\mathcal{K}_{\mathcal{X}}$ denotes the relative canonical divisor on $\mathcal{X}.$ We also assume that the corresponding complex variety $X$ is normal and thus defines a complex Fano variety. Following standard procedure we denote by $\overline{\mathcal{L}}$ a metrized line bundle, i.e. a line bundle $\mathcal{L}$ on $\mathcal{X}$ endowed with an Hermitian metric over the complex points $X$ of $\mathcal{X}.$ Arithmetic intersection numbers of metrized line bundles were introduced by Gillet-Soulé in the context of Arakelov geometry [7]. The top arithmetic intersection number of $\overline{\mathcal{L}}$ is called the height of $\overline{\mathcal{L}}.$ The height of $\overline{-\mathcal{K}_{\mathbb{P}_{\mathbb{Z}}^{n}}}$ with respect to the volume-normalized Fubini-Study metric, appearing in the previous conjecture, is explicitly given by the following formula [1, Lemma 3.6], which, essentially, goes back to [19, §5.4]:
(1.1) $(\overline{-\mathcal{K}_{\mathbb{P}_{\mathbb{Z}}^{n}}})^{n+1}=\frac{1}{2}(n+1)^{n+1}\left((n+1)\sum_{k=1}^{n}k^{-1}-n+\log(\frac{\pi^{n}}{n!})\right).$
As for the notion of K-stability it originally appeared in the context of the Yau-Tian-Donaldson conjecture for Fano manifolds $X$ (see the survey [49] for recent developments, including connections to moduli spaces and the minimal model program in birational geometry). By [29] and [31, Thm 1.6] a log Fano variety $(X,\Delta)$ over $\mathbb{C}$ is K-polystable (which is a slightly stronger condition than K-semistability) if and only if it admits a log Kähler-Einstein metric, i.e. a locally bounded metric on $X$ whose curvature current $\omega$ induces a Kähler metric with constant Ricci curvature on the complement of $\Delta$ in the regular locus of $X.$ After volume-normalization such a metric maximizes the height $(\overline{-\mathcal{K}_{(\mathcal{X},\mathcal{D})}})^{n+1}$ among all volume-normalized locally bounded metrics on $-(K_{X}+\Delta)$ with positive curvature (as shown precisely as in the case that $\mathcal{D}=0$ considered in [2, Section 2.3]).
The K-semistability of $(X,\Delta)$ implies that $(X,\Delta)$ is Kawamata Log Terminal (klt) in the usual sense of birational geometry (see Remark 2.1). An important class of klt log Fano varieties $(X,\Delta)$ is provided by (smooth) Fano orbifolds, where the coefficients of $\Delta$ are of the form $(1-1/m_{i})$ for positive integers $m_{i}.$ Diophantine aspects of Fano orbifolds have recently been explored in a number of works, building on Campana’s program [12] and its developments by Abramovich [1] (see [47] for a very recent survey). In particular, a logarithmic generalization of the Manin-Peyre conjecture for the density of rational points of bounded height on Fano varieties is proposed in [42], which, for example, is addressed for log Fano hyperplane arrangements and toric varieties in [10] and [40], respectively. See [4] for relations between height bounds, K-stability and the Manin-Peyre conjecture.
1.1. Main results
1.1.1. Toric log Fano varieties
We first consider the case when $(\mathcal{X},\mathcal{D})$ is the canonical integral model of a toric log Fano variety $(X,\Delta)$ (see [33, Section 2] and [11, Def 3.5.6]). One advantage of the logarithmic setup is that on any given toric Fano variety $X$ there exist an infinite number of toric $\mathbb{Q}-$divisors $D$ such that $-(K_{X}+\Delta)$ is a K-semistable log Fano variety. Building on [1], where the case when $D=0$ was considered, we show
Theorem 1.2.
Let $(\mathcal{X},\mathcal{D})$ be the canonical integral model of a K-semistable toric log Fano variety $(X,\Delta).$ Conjecture 1.1 holds for $(\mathcal{X},\mathcal{D})$ under anyone of the following conditions:
  • $n\leq 3$ and $X$ is $\mathbb{Q}-$factorial (equivalently, $X$ has at worst abelian quotient singularities)
  • $X$ is not Gorenstein or has some abelian quotient singularity
The starting point of the proof is the bound
(1.2) $\frac{\overline{(-\mathcal{K}_{(\mathcal{X},\mathcal{D})}})^{n+1}}{(n+1)!}\leq\frac{1}{2}\mathrm{vol}(X,\Delta)\log\left(\frac{(2\pi^{2})^{n}}{\mathrm{vol}(X,\Delta)}\right)\,\,\,\,\mathrm{vol}(X,\Delta):=\frac{-(K_{X}+\Delta)^{n}}{n!},$
shown precisely as in the case when $\Delta=0,$ considered in [1]. For $X=\mathbb{P}^{n}$ the previous theorem is verified by an explicit calculation. In the remaining case, $X\neq\mathbb{P}^{n},$ the bound in Conjecture 1.1 follows, just as in [1], from combining the bound 1.2 with the following logarithmic analog of the “gap hypothesis” introduced in [1]:
(1.3) $\text{vol}(X,\Delta)\leq\text{vol}(\mathbb{P}^{n-1}\times\mathbb{P}^{1})$
for any K-semistable $n-$dimensional Fano variety $(X,\Delta)$ such that $X\neq\mathbb{P}^{n}.$ In the case that $X$ is singular the logarithmic gap hypothesis does hold in any dimension, just as in [1]. In the non-singular case there is, for any dimension, only a finite number of toric Fano varieties $X.$ For $n\leq 6$ these appear in the database [52], which, as observed in [1], settles the gap hypothesis for $n\leq 6,$ when $\Delta=0.$ However, in the present case there is for any given toric variety $X$ an infinite number of toric divisors $\Delta$ on $X$ such that $(X,\Delta)$ is a K-semistable Fano variety. In order to establish the logarithmic gap-hypothesis 1.3 we thus introduce the following invariant of a Fano manifold $X:$
[…]
and show, by solving the corresponding optimization problem, that $S(X)\leq\text{vol}(\mathbb{P}^{n-1}\times\mathbb{P}^{1})$ when $X\neq\mathbb{P}^{n}$ and $n\leq 3.$
The invariant $S(X)$ and the corresponding maximizers $\Delta$ appear to be of independent interest in Kähler geometry. This is illustrated by some examples in Section 3.1, where we make contact with a rigidity property of the corresponding log Kähler-Einstein metric, first exhibited in [44].
1.1.2. Hyperplane arrangements
We next turn to the case when $\mathcal{X}$ is the projective space over the integers, $\mathcal{X}=\mathbb{P}_{\mathbb{Z}}^{n}$ and $\mathcal{D}$ is a hyperplane arrangement, i.e. its irreducible components are hyperplanes.
Theorem 1.3.
Conjecture 1.1 holds when $\mathcal{X}=\mathbb{P}_{\mathbb{Z}}^{n}$ and $\mathcal{D}$ is a hyperplane arrangement with simple normal crossings.
The proof employs a convexity argument to reduce the problem to the case when $\mathcal{D}$ is toric, which is covered by Theorem 3.1. The argument leverages the explicit characterization of K-semistable hyperplane arrangements established in [23] and yields the following explicit bound:
(1.4) $\frac{(\overline{-\mathcal{K}_{(\mathcal{X},\mathcal{D})}})^{n+1}}{(n+1)!}\leq\frac{1}{2}\text{vol}(X,\Delta)\log\left(\frac{(n+1)^{n}e^{2a_{n}}}{(n+1)!\text{vol}(X,\Delta)}\right),\,\,\,a_{n}=\frac{(\overline{-\mathcal{K}_{\mathbb{P}_{\mathbb{Z}}^{n}}})^{n+1}}{\left(n+1\right)^{n+1}}$
with equality iff $\mathcal{D}$ is toric.
1.1.3. Application to diagonal hypersurfaces
Given a positive integer $d$ and integers $a_{i},$ consider the diagonal hypersurface $\mathcal{X}_{a}$ of degree $d$ in $\mathbb{P}_{\mathbb{Z}}^{n+1}$ cut out by the homogeneous polynomial
$\sum_{i=0}^{n+1}a_{i}x_{0}^{d},$
The corresponding complex variety $X_{a}$ is Fano if and only if $d\leq(n+1)$ and is always K-polystable (and, in particular, K-semistable); see, for example, [51] for an algebraic proof. Using the results stated in the previous two sections we will establish Conjecture 1.1 for $\mathcal{X}_{a},$ endowed with the trivial divisor $0:$
Theorem 1.4.
Conjecture 1.1 holds for any diagonal hypersurface $\mathcal{X}_{a}$ which is Fano (i.e. $d\leq n+1)$ when the divisor $\mathcal{D}$ is trivial. More precisely,
$(\overline{-\mathcal{K}_{\mathcal{X}_{a}}})^{n+1}\leq(\overline{-\mathcal{K}_{\mathbb{P}_{\mathbb{Z}}^{n}}})^{n+1}+(1-d)(n+2-d)^{n}\sum_{i=0}^{n+1}\log|a_{i}|.$
and the inequality is strict if $d\geq 2.$
Note that the schemes $\mathcal{X}_{a}$ are mutually non-isomorphic over $\mathbb{Z},$ for any given degree $d$ of at least two. In fact, in general, they are not even isomorphic over $\mathbb{Q}.$ The proof of the previous theorem is first reduced to the case of a Fermat hypersurface, i.e. the case when $a_{i}=1.$ Expressing $\mathcal{X}_{a}$ as a Galois cover of $\mathbb{P}_{\mathbb{Z}}^{n}$ the estimate 1.4 can then be applied with $\mathcal{X}=\mathbb{P}_{\mathbb{Z}}^{n}$ and $\Delta$ the corresponding branching divisor, which reduces the problem to a simple toric case.
We recall that the Manin-Peyre conjecture has been settled for Fano hypersurfaces of sufficiently small degree [38, 39]. In particular, there is an extensive literature on the diagonal case, in connection to Waring’s classical problem, where the degree bound has been improved [41].
1.1.4. Arithmetic log surfaces
Consider a polarized arithmetic log surface $(\mathcal{X},\mathcal{D};\mathcal{L}),$ i.e. an arithmetic log surface $(\mathcal{X},\mathcal{D})$ endowed with a relatively ample line bundle $\mathcal{L}.$ Assume that the complexification $\mathcal{L}\otimes\mathbb{C}$ is isomorphic to $-K_{(X,\Delta)}$ (where $(X,\Delta)$ denotes, as before, the complexification of $(\mathcal{X},\mathcal{D})$). Given a continuous metric on $\mathcal{L}$ we define the arithmetic log Mabuchi functional of $(\mathcal{X},\mathcal{D};\overline{\mathcal{L}})$ by
(1.5) $\mathcal{M}_{(\mathcal{X},\mathcal{D})}(\overline{\mathcal{L}}):=\frac{1}{2}\overline{\mathcal{L}}^{2}+\overline{\mathcal{K}}_{(\mathcal{X},\mathcal{D})}\cdot\overline{\mathcal{L}},$
where $\overline{\mathcal{K}}_{(\mathcal{X},\mathcal{D})}$ is endowed with the volume-normalized metric induced by the curvature current $\omega$ of $\overline{\mathcal{L}},$ assuming that this metric on $\overline{\mathcal{K}}_{(\mathcal{X},\mathcal{D})}$ is continuous. When $\mathcal{D}=0$ the functional $\mathcal{M}_{(\mathcal{X},\mathcal{D})}(\overline{\mathcal{L}})$ coincides with the arithmetic Mabuchi functional introduced in [37], up to normalization (see [2, Section 6.4]). For a given integral model $\mathcal{M}_{(\mathcal{X},\mathcal{D})}(\overline{\mathcal{L}})$ is minimized on a log Kähler-Einstein metric, if such a metric exists, and then
$\mathcal{M}_{(\mathcal{X},\mathcal{D})}(\overline{\mathcal{L}})=-(\overline{-\mathcal{K}_{(\mathcal{X},\mathcal{D})}})^{2}/2,$
if the metric is volume-normalized and $\mathcal{L}=-\overline{\mathcal{K}}_{(\mathcal{X},\mathcal{D})}.$
Theorem 1.5.
Let $(\mathcal{X},\mathcal{D};\mathcal{L})$ be a polarized arithmetic log surface $(\mathcal{X},\mathcal{D};\mathcal{L})$ such that the complexification $(X,\Delta)$ of $(\mathcal{X},\mathcal{D})$ is a K-semistable Fano variety and $\mathcal{L}\otimes\mathbb{C}=-K_{(X,\Delta)}.$ Then
$\mathcal{M}_{(\mathcal{X},\mathcal{D})}(\overline{\mathcal{L}})\geq\mathcal{M}_{(\mathbb{P}_{\mathbb{Z}}^{1},0)}(\overline{-\mathcal{K}_{\mathbb{P}_{\mathbb{Z}}^{1}}})\,\,\,\left(=-1-\log\pi\right)$
where $-\mathcal{K}_{\mathbb{P}_{\mathbb{Z}}^{1}}$ is endowed with the Fubini-Study metric. Moreover, equality holds iff $(\mathcal{X},\mathcal{D})$ is isomorphic to $(\mathbb{P}_{\mathbb{Z}}^{1},0)$ and $\mathcal{L}$ is isomorphic to $-\mathcal{K}_{\mathbb{P}_{\mathbb{Z}}^{1}},$ endowed with a metric coinciding with the Fubini-Study metric, up to the application of an automorphism of $\mathbb{P}_{\mathbb{Z}}^{1}$ and a scaling of the metric.
Corollary 1.6.
Conjecture 1.1 holds for arithmetic log Fano surfaces.
In the setup of the previous theorem the corresponding complex variety $X$ is always equal to $\mathbb{P}^{1}$ and thus $(X,\Delta)$ is a hyperplane arrangement. Accordingly, applying Theorem 1.3, the proof of Theorem 1.5 is reduced to showing that the canonical integral model $(\mathcal{X}_{c},\mathcal{D}_{c};-\mathcal{K}_{(\mathcal{X}_{c},\mathcal{D}_{c})})$ of $(X,\Delta;-K_{(X,\Delta)})$ obtained by setting $\mathcal{X}_{c}=\mathbb{P}_{\mathbb{Z}}^{1}$ and taking $\mathcal{D}_{c}$ to be the Zariski closure of $\Delta_{\mathbb{Q}}$ in $\mathcal{X}_{c}$ minimizes $\mathcal{M}_{(\mathcal{X},\mathcal{D})}(\overline{\mathcal{L}})$ over all integral models $(\mathcal{X},\mathcal{D};\mathcal{L})$ of $(X,\Delta;-K_{(X,\Delta)}$), for any fixed metric on $-K_{(X,\Delta)}.$ This minimization property can be viewed as logarithmic version of Odaka’s minimization conjecture (proposed in any dimension in [37]). Our proof builds on [37], leveraging log canonical thresholds. We also show that the minimum is uniquely attained for $(\mathbb{P}_{\mathbb{Z}}^{1},0).$ See also [21] for very recent progress on Odaka’s minimization conjecture in another direction.
1.2. Outlook on the case of $\mathbb{P}_{\mathbb{Z}}^{1}$ endowed with a divisor with three components
Consider now the special case of $(\mathbb{P}^{1},\mathcal{D}_{c}),$ where $\mathcal{D}_{c}$ is the canonical model of a divisor $\Delta_{\mathbb{Q}}$ on $\mathbb{P}^{1}$ with three components and $-K_{(\mathbb{P}^{1},\Delta_{\mathbb{Q}})}$ is endowed with the volume-normalized log Kähler-Einstein metric. In this case an explicit formula for the canonical height $\overline{(-\mathcal{K}_{(\mathbb{P}_{\mathbb{Z}}^{1},\mathcal{D}_{c})}})^{2}$ is established in the sequel [3] to the present work. More precisely, expressing
$\Delta_{\mathbb{Q}}=\sum_{i=1}^{3}w_{i}p_{i},\,\,\,V=2-\sum_{i=1}^{3}w_{i}>0$
(where $V$ is the algebraic degree of $-K_{(\mathbb{P}^{1},\Delta_{\mathbb{Q}})})$ it is shown in [3] that, when $(\mathbb{P}_{1},\Delta_{\mathbb{Q}})$ is K-polystable,
(1.6) $\frac{(\overline{-\mathcal{K}_{(\mathbb{P}_{\mathbb{Z}}^{1},\mathcal{D}_{c})}})^{2}}{2V}=\frac{1}{2}(1+\log\pi-\log\frac{V}{2})-\frac{\gamma(0,\frac{V}{2})+\sum_{i=1}^{3}\gamma(w_{i},w_{i}+\frac{V}{2})}{V},$
where $\gamma(a,b)$ may be expressed in terms of Hurwitz zeta function $\zeta(s,x)$ and its derivative $\zeta^{\prime}(s,x)$ wrt $s:$
$\gamma(a,b)=F(b)+F(1-b)-F(a)-F(1-a),\,\,\,F(x):=\zeta(-1,x)+\zeta^{\prime}(-1,x).$
Moreover, the right hand side of formula 1.6 is shown to extend real-analytically, wrt the coefficients $w_{i},$ to the case $K_{(\mathbb{P}^{1},\Delta_{\mathbb{Q}})}>0$ (i.e. $V<0),$ as long as $w_{i}\in]0,1].$ In this case formula 1.6 computes $(\mathcal{K}_{(\mathbb{P}_{\mathbb{Z}}^{1},\mathcal{D}_{c})})^{2}/2V$ and can be related to Kudla’s program [28] and the Maillot-Rössler conjectures [34, 35], expressing the height of Shimura varieties $(\mathcal{X},\mathcal{D}),$ wrt canonical metrics, in terms of the Dedekind zeta function $\zeta_{\mathbb{F}}(s)$ of the number field $\mathbb{F}$ attached to $(\mathcal{X},\mathcal{D})$ and its derivative $\zeta^{\prime}_{\mathbb{F}}(s)$ at $s=-1.$ More precisely, this connection concerns the $18$ Shimura varieties that are isomorphic to $(\mathbb{P}^{1},\Delta)$ over $\mathbb{F},$ for a particular orbifold divisor $\Delta.$ In particular, combining formula 1.6 with the recent work [50] on modular heights yields information about the canonical integral models of the corresponding Shimura curves.
1.3. Acknowledgments
Thanks to Dennis Eriksson, Mattias Jonsson, Gerard Freixas i Montplet and Yuji Odaka for discussions. This work was supported by a Wallenberg Scholar grant from the Knut and Alice Wallenberg foundation.
2. General setup
2.1. Log Fano varieties over $\mathbb{C}$ and volume-normalized metrics on $-(K_{X}+\Delta)$
[…]
2.1.1. Local representations of metrics and measures
[…]
(2.1) $\phi_{U}:=-\log(\left\|e_{U}\right\|^{2}).$
The curvature current of the metric may then, locally, be expressed as
$dd^{c}\phi_{U}:=\frac{i}{2\pi}\partial\bar{\partial}\phi_{U}$
Accordingly, as is customary, we will symbolically denote by $\phi$ a given continuous Hermitian metric on $L$ and by $dd^{c}\phi$ its curvature current. We will denote by $\mathcal{C}^{0}(L)\cap\text{PSH$(L)$}$ the space of all continuous metrics on $L$ whose curvature current is positive, $dd^{c}\phi\geq 0$ (which means that $\phi_{U}$ is plurisubharmonic, or psh, for short).
Given a log Fano pair $(X,\Delta)$ the measure corresponding to a given continuous metric $\phi$ on $-K_{(X,\Delta)}$ may be locally on $X_{reg}$ be expressed as
$\mu_{\phi}=e^{-\phi_{U}}\left|s_{U}\right|^{-2}(\frac{i}{2})^{n^{2}}dz\wedge d\bar{z},\,\,\,dz:=dz_{1}\wedge\cdots\wedge dz_{n}$
by taking $e_{U}=\partial/\partial z_{1}\wedge\cdots\wedge\partial/\partial z_{n}\otimes e_{\Delta}$ where $e_{\Delta}$ is a local trivialization of the $\mathbb{Q}-$line bundle over $X_{reg}$ corresponding to the divisor $\Delta$ and $s_{U}e_{\Delta}$ is the (multi-valued) holomorphic section cutting out $\Delta.$
2.1.2. Log Kähler-Einstein metrics
Given a log Fano pair $(X,\Delta)$ a metric $\phi$ on $(X,\Delta)$ is said to be a log Kähler-Einstein metric, if $\phi$ is a locally bounded metric and its curvature current $dd^{c}\phi$ induces a Kähler metric with constant positive Ricci curvature on the complement of $\Delta$ in $X_{reg}$ [5]. When $X$ is smooth any log Kähler-Einstein metric is, in fact, continuous (see [25, 20] for more general higher order regularity results).
2.1.3. K-semistability
We next recall the definition of K-semistability in terms of intersection numbers (see the survey [49] for more background). A test configuration for a log Fano pair $(X,\Delta)$ is a $\mathbb{C}^{*}-$equivariant normal model $(\mathscr{X},\mathscr{L})$ for $(X,-(K_{(X,\Delta)})$ over the complex affine line $\mathbb{C}.$ More precisely, $\mathscr{X}$ is a normal complex variety endowed with a $\mathbb{C}^{*}-$action $\rho$, a $\mathbb{C}^{*}-$equivariant holomorphic projection $\pi$ to $\mathbb{C}$ and a relatively ample $\mathbb{C}^{*}-$equivariant $\mathbb{Q}-$line bundle $\mathscr{L}$ (endowed with a lift of $\rho):$
(2.2) $\pi:\mathcal{\mathscr{X}}\rightarrow\mathbb{C},\,\,\,\,\,\mathscr{L}\rightarrow\mathscr{X},\,\,\,\,\,\,\rho:\,\,\mathscr{X}\times\mathbb{C}^{*}\rightarrow\mathscr{X}$
such that the fiber of $\mathscr{X}$ over $1\in\mathbb{C}$ is equal to $(X,-(K_{(X,\Delta)}).$ A log Fano pair $(X,\Delta)$ is said to be K-semistable if the Donaldson-Futaki invariants $\text{DF}_{\Delta}(\mathscr{X},\mathscr{L})$ are non-negative for any test configuration $(\mathscr{X},\mathscr{L})$ of $(X,\Delta):$
$n!\text{DF}_{\Delta}(\mathscr{X},\mathscr{L})=\frac{n}{(n+1)}\overline{\mathscr{L}}^{n+1}+\mathscr{K}_{(\mathcal{\mathscr{\overline{X}}},\mathscr{D})/\mathbb{P}^{1}}\cdot\mathcal{\overline{\mathscr{L}}}^{n},$
where $\overline{\mathscr{L}}$ denotes the $\mathbb{C}^{*}-$equivariant extension of $\mathscr{L}$ to the $\mathbb{C}^{*}-$equivariant compactification $\mathscr{\overline{X}}$ of $\mathscr{X}$ over $\mathbb{P}^{1}$ and $\mathscr{K}_{(\mathcal{\mathscr{\overline{X}}},\mathscr{D})/\mathbb{P}^{1}}$ denotes the relative log canonical divisor of the pair $(\mathscr{\overline{X}}$,$\mathscr{D}$) with $\mathscr{D}$ denoting the flat closure in $\mathscr{\overline{X}}$ of the $\mathbb{C}^{*}-$orbit of the divisor $\Delta.$
Remark 2.1.
If a log Fano variety $(X,\Delta)$ is K-semistable, then $(X,\Delta)$ is klt [9, Cor 9.6]. When $X$ is non-singular and $\Delta$ has simple normal crossings this means that all the coefficients of $\Delta$ along its irreducible components are strictly smaller than $1$
2.2. Arithmetic log Fano varieties
[…]
$\pi:\,\mathcal{X}\rightarrow\text{Spec $\mathbb{Z}$.}$.
A log pair $(\mathcal{X},\mathcal{D})$ over $\mathbb{Z}$ (also called an arithmetic log variety) of relative dimension $n$ is an arithmetic variety $\mathcal{X}$ endowed with an effective $\mathbb{Q}-$divisor $\mathcal{D}$ on $\mathcal{X}$ such that $\mathcal{K}_{\mathcal{X}}+\mathcal{D}$ is $\mathbb{Q}-$Cartier, i.e. defines a $\mathbb{Q}-$line bundle, where $\mathcal{K}_{\mathcal{X}}$ denotes the relative canonical divisor on $\mathcal{X}$ (see [27, Section 1.1]). The complexification of $(\mathcal{X},\mathcal{D})$ will be denoted by $(X,\Delta)$ and $(\mathcal{X},\mathcal{D})$ will be called an integral model of $(X,\Delta)$ (although, strictly speaking, $(\mathcal{X},\mathcal{D})$ is an integral model of the corresponding log pair over $\mathbb{Q}$). A log pair $(\mathcal{X},\mathcal{D})$ over $\mathbb{Z}$ will be called an arithmetic log Fano variety if $-(\mathcal{K}_{\mathcal{X}}+\mathcal{D})$ is relatively ample and the corresponding complex variety $X$ is normal. In particular, $(X,\Delta)$ is a log Fano variety over $\mathbb{C}.$
More generally, an arithmetic variety $\mathcal{X}$ endowed with a relatively ample line bundle $\mathcal{L}$ will be said to be polarized.
2.3. Arithmetic intersection numbers and heights
We recall some well-known facts about heights (see[2] for more background and references). A metrized line bundle $\overline{\mathcal{L}}$ is a line bundle $\mathcal{L}\rightarrow\mathcal{X}$ such that the corresponding line bundle $L\rightarrow X$ is endowed with a metric, that we shall denote by $\phi$ (as in Section 2.1.1); ¯$\overline{\mathcal{L}}:=\left(\mathcal{L},\phi\right).$ The $\chi-$arithmetic volume of a polarized arithmetic variety $(\mathcal{X},\mathcal{L})$ is defined by
[…]
where $H^{0}(\mathcal{X},k\mathcal{L})\otimes\mathbb{R}$ may be identified with the subspace of real sections in $H^{0}(X,kL).$ More generally, $\widehat{\text{vol}}_{\chi}\left(\overline{\mathcal{L}}\right)$ is naturally defined for $\mathbb{Q}-$line bundles, since it is homogeneous with respect to tensor products of $\overline{\mathcal{L}}:$
(2.4) $\widehat{\text{vol}}_{\chi}\left(m\overline{\mathcal{L}}\right)=m^{n+1}\widehat{\text{vol}}_{\chi}\left(\overline{\mathcal{L}}\right),\,\,\,\text{if $m\in\mathbb{Z}_{+}$}$
Moreover, $\widehat{\text{vol}}_{\chi}\left(\overline{\mathcal{L}}\right)$ is additively equivariant with respect to scalings of the metric:
(2.5) $\widehat{\text{vol}}_{\chi}\left(\mathcal{L},\phi+\lambda\right)=\widehat{\text{vol}}_{\chi}\left(\overline{\mathcal{L}}\right)+\frac{\lambda}{2}\text{vol$(L)$},\,\,\,\text{if }\lambda\in\mathbb{R}.$
If the metric on $L$ has positive curvature current (i.e. if $\phi$ is psh), then, by the arithmetic Hilbert-Samuel theorem,
(2.6) $\widehat{\text{vol}}_{\chi}\left(\overline{\mathcal{L}}\right)=\frac{\overline{\mathcal{L}}^{n+1}}{(n+1)!},$
where ¯$\overline{\mathcal{L}}^{n+1}$ denotes the top arithmetic intersection number in the sense of Gillet-Soulé [19], which, defines the height of $\mathcal{X}$ with respect to $\overline{\mathcal{L}}$ [18, 7]. For the purpose of the present paper formula 2.6 may be taken as the definition of ¯$\overline{\mathcal{L}}^{n+1}$ (arithmetic intersections between general $n+1$ metrized line bundles could then be defined by polarization, i.e. using multilinearity). Following standard practice we will use the shorthand $h_{\phi}(\mathcal{X},\mathcal{L})$ for the height $(\mathcal{L},\phi)^{n+1}$ and $\hat{h}_{\phi}(\mathcal{X},\mathcal{L})$ for the normalized height:
$h_{\phi}(\mathcal{X},\mathcal{L}):=(\mathcal{L},\phi)^{n+1},\,\,\,\hat{h}_{\phi}(\mathcal{X},\mathcal{L}):=\frac{(\mathcal{L},\phi)^{n+1}}{(n+1)L^{n}}.$
The definition of $\hat{h}_{\phi}(\mathcal{X},\mathcal{L})$ is made so that
$\hat{h}_{\phi+\lambda}(\mathcal{X},\mathcal{L})=\hat{h}_{\phi}(\mathcal{X},\mathcal{L})+\lambda/2,\,\,\,\text{if }\lambda\in\mathbb{R}$
We also recall that, given two continuous psh metrics $\phi$ and $\phi_{0}$ on the complexification $L\rightarrow X$ of $\mathcal{L}\rightarrow\mathcal{X},$ we have that
(2.7) $2h\left(\mathcal{L},\phi\right)-2h\left(\mathcal{L},\phi_{0}\right)=\mathcal{E}(\phi,\phi_{0}):=\frac{1}{(n+1)!}\int_{X}(\phi-\phi_{0})\sum_{j=0}^{n}(dd^{c}\phi)^{j}\wedge(dd^{c}\phi_{0})^{n-j}.$
2.4. The canonical height of an arithmetic log Fano variety
We define the canonical height $h_{\text{can}}(\mathcal{X},\mathcal{D})$ of an arithmetic log Fano variety $(\mathcal{X},\mathcal{D})$ by
$h_{\text{can}}(\mathcal{X},\mathcal{D}):=\sup\left\{h_{\phi}\left(-\mathcal{K}_{(\mathcal{X},\mathcal{D})}\right):\,\text{$\phi\,\,\text{$\text{cont. }$psh,}$}\int_{X}\mu_{\phi}=1\right\}$
(when $\mathcal{D}=0$ we shall use the short hand $h_{\text{can}}(\mathcal{X})$ for $h_{\text{can}}(\mathcal{X},0)$). As shown precisely as in the case $\mathcal{D}=0,$ considered in [2], $h_{\text{can}}(\mathcal{X},\mathcal{D})<\infty$ iff the corresponding log Fano variety $(X,\Delta)$ over $\mathbb{C}$ is K-semistable. Moreover, $(X,\Delta)$ is K-polystable iff the sup defining $h_{\text{can}}(\mathcal{X},\mathcal{D})$ is attained at some continuous metric $\phi,$ namely a log Kähler-Einstein metric. Hence, if $(X,\Delta)$ is K-polystable, then the canonical height $h_{\text{can}}(\mathcal{X},\mathcal{D})$ is computed by any volume-normalized log Kähler-Einstein metric.
3. Toric log Fano varieties
A log pair $(X,D)$ over $\mathbb{C}$ is said to be toric if $X$ and $D$ are toric, i.e. if $X$ is toric and the $\mathbb{Q}-$divisor $D$ is invariant under the torus action on $X.$ Any toric log Fano variety admits a canonical integral model $(\mathcal{X},\mathcal{D})$ which is log Fano (see [33, Section 2] and [11, Def 3.5.6]). In this section we will prove the following
Theorem 3.1.
Let $(\mathcal{X},\mathcal{D})$ be the canonical integral model of a K-semistable toric log Fano variety $(X,D).$ Conjecture 1.1 holds for $(\mathcal{X},\mathcal{D})$ under anyone of the following conditions:
  • $n\leq 3$ and $X$ is $\mathbb{Q}-$factorial (equivalently, $X$ has at worst abelian quotient singularities)
  • $X$ is not Gorenstein or has some abelian quotient singularity
We start by introducing some notation, following [6]. Given a toric log Fano variety $(X,D)$ set $L=-(K_{X}+\Delta)$ and denote by $P$ the corresponding moment polytope in $\mathbb{R}^{n}.$ Then
(3.1) $P=\left\{p\in\mathbb{R}^{n}:\,\,\left\langle l_{F},p\right\rangle\geq-a_{F},\,\,\forall F\right\},$
where $a_{F}\in]0,1]$ (generalizing the Fano case when $a_{F}=1\forall F$; see [6]) and $l_{F}$ is a primitive integer vector. As shown in [6] $(X,\Delta)$ is K-semistable iff 00 is the barycenter of $P$ iff the log Ding functional $\mathcal{D}_{\psi_{P}}$ is bounded from below. Moreover, the infimum of $\mathcal{D}_{\psi_{P}}$ is attained at a $T-$invariant psh metric $\phi$ on $L.$ We will identify the metric $\phi$ with a continuous convex function on $\mathbb{R}^{n}$ as in [2]. More precisely, on (*$(\mathbb{C}^{*})^{n}\hookrightarrow X$, let $x_{i}=\log(|z_{i}|^{2})$. Trivializing $-(K_{X}+\Delta)$ with $\frac{\mathrm{d}z_{1}}{z_{1}}\wedge...\wedge\frac{\mathrm{d}z_{n}}{z_{n}}\otimes s_{U}e_{\Delta}$ over $U=(\mathbb{C}^{*})^{n}$, and abusing notation slightly, we let $\phi(x):=\phi_{U}(z)$ in the chosen trivialization over $U=(\mathbb{C}^{*})^{n}$. Then $\phi$ as a function of $x$ is a continuous convex function on $\mathbb{R}^{n}$ and as in formula 3.8 in [2], we still have that
[…]
(since the support of $\mathcal{D}$ is contained in the complement of (*$(\mathbb{C}^{*})^{n}$ in $X).$ Thus the inequality in Proposition 3.7 in [2] generalizes to the canonical toric model $\mathcal{L}$ of $L$ (which coincides with $-\mathcal{K}_{(\mathcal{X},\mathcal{D})}):$
(3.2) $2\widehat{\mathrm{vol}}_{\chi}\left(-\mathcal{K}_{(\mathcal{X},\mathcal{D})},\phi\right)\leq-\mathrm{vol}(X,\Delta)\log\left(\frac{\mathrm{vol}(X,\Delta)}{(2\pi^{2})^{n}}\right)\,\,\mathrm{vol}(X,\Delta):=\mathrm{vol}(-K_{(X,\Delta}).$
We will first prove Theorem 3.1 in the case that $X=\mathbb{P}^{n},$ using the following lemma, formulated in terms of the divisor $D_{0}$ cut out by the $T_{c}-$invariant element of $H^{0}(X,-K_{X})$ (given by $\frac{\mathrm{d}z_{1}}{z_{1}}\wedge...\wedge\frac{\mathrm{d}z_{n}}{z_{n}}$ over (*$(\mathbb{C}^{*})^{n}$). In other words,
$D_{0}=\sum_{F}D_{F},$
where $D_{F}$ is the irreducible divisor corresponding to the facet $F$ of the moment polytope corresponding to $X$ (see [6]). The lemma is a special case of formula Proposition 3.12 from [2].
Lemma 3.2.
Let $\mathcal{X}$ be the canonical integral model of an $n-$dimensional K-semistable toric Fano variety $X$ and denote by $D_{0}$ the standard anti-canonical divisor on $X.$ Then
$\frac{(\overline{-\mathcal{K}_{(\mathcal{X},(1-t)\mathcal{D}_{0})}})^{n+1}/(n+1)!}{\left(-(K_{X}+(1-t)D_{0})\right)^{n}/n!}=\frac{(\overline{-\mathcal{K}_{\mathcal{X}}})^{n+1}/(n+1)!}{\left(-K_{X}\right)^{n}/n!}-\frac{1}{2}\log(t^{n})\,\,\,\,\,\,\,t^{n}=\left(\frac{\left(-(K_{X}+(1-t)D_{0})\right)^{n}}{\left(-K_{X}\right)^{n}}\right)$
with respect to the volume-normalized Kähler-Einstein metrics.
We next deduce the following
Lemma 3.3.
Let $(\mathcal{X},\mathcal{D})$ be a toric K-semistable log Fano variety such that $\mathcal{X}=\mathbb{P}_{\mathbb{Z}}^{n}.$ Then $(\overline{-\mathcal{K}_{(\mathcal{X},\mathcal{D})}})^{n+1}\leq(\overline{-\mathcal{K}_{\mathbb{P}_{\mathbb{Z}}^{n}}})^{n+1}$ with equality iff $\mathcal{D}=0.$
Proof.
First observe that there exists $t\in[0,1]$ such that $\mathcal{D}=(1-t)\mathcal{D}_{0}=:\mathcal{D}_{t}.$ This is a special case of [23, Cor 1.6], which applies to $\mathbb{P}^{n},$ in any dimension $n,$ using that toric log Fano varieties are never uniformly K-stable. It will thus be enough to show that $t\mapsto(\overline{-\mathcal{K}_{(\mathbb{P}_{\mathbb{Z}}^{n},\mathcal{D}_{t})}})^{n+1}$ is increasing on $[0,1]$ (and thus its maximum is attained at $t=1).$ By the previous lemma
$\frac{2(\overline{-\mathcal{K}_{(\mathbb{P}_{\mathbb{Z}}^{n},\mathcal{D}_{t})}})^{n+1}/(n+1)!}{(-K_{\mathbb{P}^{n}})^{n}/n!}=t^{n}2\frac{(\overline{-\mathcal{K}_{\mathbb{P}_{\mathbb{Z}}^{n}}})^{n+1}/(n+1)!}{(-K_{\mathbb{P}^{n}})^{n}/n!}-t^{n}\log(t^{n}).$
Differentiating wrt $(t^{n})$ reveals that the right hand side above is increasing with respect to $t$ iff $2\frac{(\overline{-\mathcal{K}_{\mathbb{P}_{\mathbb{Z}}^{n}}})^{n+1}/(n+1)!}{(-K_{\mathbb{P}^{n}})^{n}/n!}\geq 1$ The latter inequality is indeed satisfied, as follows from the explicit formula 1.1. ∎
Combining the universal bound 3.2 with Lemma 3.8 from [2], all that remains to prove Theorem 3.1 is to establish the “logarithmic gap hypothesis”
(3.3) $\text{vol}(X,\Delta)\leq\text{vol}(\mathbb{P}^{n-1}\times\mathbb{P}^{1}).$
Proposition 3.4.
The logarithmic gap hypothesis holds for all toric K-semistable log Fano varieties (manifolds) $(X,\Delta)$ such that $X\neq\mathbb{P}^{n}$ iff the following bound holds for all Fano varieties (manifolds) $X\neq\mathbb{P}^{n}$
[…]
The “logarithmic gap hypothesis” holds for all log Fano varieties $(X,\Delta)$ such that $X$ is $\mathbb{Q}-$factorial and of dimension $n\leq 3$ and for any dimensions $n$ if $X$ has some abelian quotient singularity or if $X$ is not Gorenstein.
Proof.
Since, trivially, $\text{vol}(X,\Delta)\leq S(X)$ the first equivalence follows directly from the definitions. Next, let us show the last statement of the proposition, first assuming that $X$ is singular, which means that the moment polytope $P$ of $(X,\Delta)$ is “singular” in the sense that there exists a vertex of $\partial P$ such that the corresponding primitive vectors $l_{F_{1}},...,l_{F_{n}}$ do not generate $\mathbb{Z}^{n}.$ It follows from the proof of Lemma 3.9 from [2] that
$\mathrm{vol}(P)\leq\frac{1}{2}(n+1)^{n}/n!\leq\text{vol}(\mathbb{P}^{n-1}\times\mathbb{P}^{1}))$
Indeed, since $a_{F}\leq 1$ the first inequality follows from the inequality (3.13) from [2], using that $\delta\geq 2,$ according to the singularity assumption on $P$ (for the the second inequality see formula (3.14) from [2]). All that remains is thus to show the bound 3.4 for $S(X)$ when $n\leq 3$ and $X$ is non-singular. First assume that $n=2.$ This means, by classical classification results, that $X$ is either $\mathbb{P}^{1}\times\mathbb{P}^{1}$ or the blow-up $X^{(m)}$ of $\mathbb{P}^{2}$ in $m$ points for $m\leq 3.$ But $(-K_{X^{(m)}})^{2}=(-K_{\mathbb{P}^{2}})^{2}-m$ and thus $\text{vol}(X_{1})\leq 4=\text{vol}(\mathbb{P}^{n-1}\times\mathbb{P}^{1}),$ proving the bound 3.4. Finally, consider the case when $n=3.$ Starting with the trivial bound $\text{vol}(X,\Delta)\leq\text{vol}(X)$ it follows the classification [52] of all non-singular toric Fano varieties of dimension 3 that it is enough to show that the bound 3.4 holds when $X$ is $\mathbb{P}^{3}$ blown-up in one point or $\mathbb{P}(\mathcal{O}(1)\oplus\mathcal{O}(2))$ (whose degrees are $56$ and $62,$ respectively). According to the following proposition the corresponding invariants $S(X)n!$ are, approximately, given by $41.8$ and $30.3$, respectively, which are well below the degree $54$ of $\mathbb{P}^{2}\times\mathbb{P}^{1},$ as desired. ∎
3.1. The invariant $S(X)$ for $n\leq 3$
In the proof above we used the following result.
Proposition 3.5.
After rounding to the nearest decimal place the invariant $n!S(X)$ (formula 3.4) is given by $41.8$ and $30.3$ when $X$ equals $\mathbb{P}^{3}$ blown up in one point and $\mathbb{P}(\mathcal{O}\oplus\mathcal{O}(2)),$ respectively.
Proof.
Given a convex subset $P$ of $\mathbb{R}^{n}$ let
$s(P):=\sup\left\{\text{vol}(P_{0}):\,P_{0}\subset P,\,b_{P_{0}}=0\right\},$
where $P_{0}$ is a closed subset of $P$ with barycenter $b_{P_{0}}$ at the origin. We will compute $s(P)$ when $P$ is the moment polytope of the manifolds $X$ appearing in the proposition, showing at the same time that $s(P)=S(X).$ The moment polytopes $P$ of both $\mathbb{P}^{3}$ blown up in one point and $\mathbb{P}(\mathcal{O}\oplus\mathcal{O}(2))$ are of the form a simplex, with a simplex subset removed, by chopping off a vertex (see ID 20 and ID7 in the database [52])). After a general linear transformation, they are of the form $(a\Delta_{3}-\mathbf{1})-(b\Delta_{3}-\mathbf{1})$ where $\Delta_{3}$ is the standard unit simplex in dimension three, $\mathbf{1}$ is the vector with all ones and $a$ and $b$ are positive real numbers. For $\mathbb{P}^{3}$ blown up in one point we can transform the moment polytope to $(4\Delta_{3}-\mathbf{1})\backslash(2\Delta_{3}-\mathbf{1})$ and for $\mathbb{P}(\mathcal{O}\oplus\mathcal{O}(2))$ we get $(5\Delta_{3}-\mathbf{1})\backslash(\Delta_{3}-\mathbf{1})$. In the first case, the linear transformation is unimodular, but in the second case the transformation has determinant $2$ This will not matter when computing $s(P)$ as long as we correct for the non-unit determinant. Next we compute the barycenter $b_{P}$ of these polytopes, a simple task using the explicit barycenter of the standard unit simplex, $b_{\Delta_{n}}=\mathbf{1}/(n+1),$ and then scaling and linearity properties of the volume times the barycenter. The barycenter of $(a\Delta_{3}-\mathbf{1})\backslash(b\Delta_{3}-\mathbf{1})$ is given by $\frac{a^{3}/3!(a/4-1)-b^{3}/3!(b/4-1)}{a^{3}/3!-b^{3}/3!}\mathbf{1}$. Next we use a general fact, to be proved in the lemma below, stating that the closed subset $P^{\prime}$ of $P$ which maximizes volume, with the relaxed constraint
(3.5) $b_{P^{\prime}}\cdot\mathbf{1}=0$
is the one given by $P\cap H$ where $H$ is a half-space with normal $\mathbf{1}$. In our case, by symmetry, this $P^{\prime}$ automatically satisfies the stronger constraint $b_{P^{\prime}}=0.$ Moreover, since the boundary of $P\cap H$ is parallel to a facet of $P$ it corresponds to a divisor $\Delta$ on $X$ defining a log Fano pair $(X,\Delta).$ Thus $(X,\Delta)$ is also the K-semistable log Fano pair realizing the sup in the definition of $S(X),$ showing that $s(P)=S(X).$ We can find $H$ by imposing the constraint. We introduce the weight $w$ such that
$P\cap H=((a-w)\Delta_{3}-\mathbf{1})\backslash(b\Delta_{3}-\mathbf{1}).$
From here it is clear that if $b_{P^{\prime}}\cdot\mathbf{1}=0,$ then, in fact, the entire barycenter will vanish and the condition $b_{P^{\prime}}\cdot\mathbf{1}=0$ turns into the following fourth order polynomial equation for $w:$
$(a-w)^{3}/3!((a-w)/4-1)-b^{3}/3!(b/4-1)=0.$
[…]
In the above proof we used the following
Lemma 3.6.
Let $P$ be a closed subset of $\mathbb{R}^{n}$ with the origin as an interior point. Given $v\in\mathbb{R}^{n}$ assume that $\int_{P}x\cdot v>0$. Then the maximum
$\max_{Q\subset P:\int_{Q}v\cdot x\mathrm{d\lambda(x)=0}}\int_{Q}\mathrm{d}\lambda$
is attained at $Q=P\cap H$ with $H$ a closed half-space with outward pointing normal $v$. Here $\mathrm{d}\lambda$ is Lebesgue measure.
Proof.
Without loss of generality we can assume that $v=(0,...,0,1).$ Denote by $(x_{1},x_{2},...,x_{n-1},y)$ the coordinates on $\mathbb{R}^{n}.$ Since the origin is an interior point of $P$ and $\int_{P}x\cdot v>0$ there is a closed half-space $H$ as in the lemma satisfying $\int_{P\cap H}y\mathrm{d\lambda=0}$. Hence, any candidate $Q$ for the maximum in question satisfies $\int_{P\cap H}y\mathrm{d\lambda=\int_{Q}y\mathrm{d\lambda}.}$ Subtracting the left hand side from the right hand side and vice versa yields $\int_{P\cap H\backslash Q}y\mathrm{d\lambda=\int_{Q\backslash P\cap H}y\mathrm{d\lambda}.}$ Since $\sup_{P\cap H\backslash Q}y\leq\inf_{Q\backslash(P\cap H)}y$ it follows that vol$(P\cap H\backslash Q)\geq\text{vol($Q\backslash P\cap H)$}$ which, in turn, implies that vol$(P\cap H)\geq\text{vol($Q)$}$, as desired. ∎
In fact, with just a slight variation of the argument above, any maximizer must be of the special form above and, in addition, assuming connectedness of $P$, the maximizer is unique. The proof of the previous proposition thus reveals that the unique toric divisor $\Delta$ on $X$ realizing the sup defining the invariant $S(X)$ is a multiple of the prime divisor $D_{F}$ defined by the zero-section of $\mathbb{P}(\mathcal{O}\oplus\mathcal{O}(2))\rightarrow\mathbb{P}^{2}$ and hyperplane “at infinity” in $\mathbb{P}^{3}$ blown up at the origin in $\mathbb{C}^{3}\subset\mathbb{P}^{3},$ respectively (i.e the zero-section of $\mathbb{P}(\mathcal{O}\oplus\mathcal{O}(1))\rightarrow\mathbb{P}^{2}).$ A similar argument also applies when $X$ is the blow-up of $\mathbb{P}^{2}$ at the origin in $\mathbb{C}^{2}$ (i.e. the first Hirzebruch surface $\mathbb{P}(\mathcal{O}\oplus\mathcal{O}(1))\rightarrow\mathbb{P}^{2}).$ The unique maximizer for the invariant $S(X)$ is then a log Fano pair $(X,\Delta)$ for a multiple of the hyperplane $D$ “at infinity” (i.e. the zero-section of $\mathbb{P}(\mathcal{O}\oplus\mathcal{O}(1))\rightarrow\mathbb{P}^{2}).$ Interestingly, this K-polystable log pair $(X,\Delta)$ was also singled out in [44, Cor 1.5] by the following rigidity property (answering a question of Cheltsov): it admits a rigid Kähler-Einstein metric in the sense that for any other multiple $cD$ the log pair $(X,cD)$ does not admit a Kähler-Einstein metric. The same rigidity property holds for the two three-dimensional log pairs discussed above (since there is a unique half-space $H$ satisfying the constraint in formula 3.5).
3.2. Estimates on the canonical height
Theorem 1.3 from [2] (and its corollary) generalizes directly to the case of log Fano pairs and their Kähler-Einstein metrics in any relative dimension $n$ (with the same proof, by letting $P$ be the moment polytope corresponding to $(X,\Delta)):$
$\frac{1}{2}\mathrm{vol}(X,\Delta)\log\left(\frac{n!m_{n}\pi^{n}}{\mathrm{vol}(X,\Delta)}\right)\leq\frac{h_{\text{can}}(\mathcal{X},\mathcal{D})}{(n+1)!}\leq\frac{1}{2}\mathrm{vol}(X,\Delta)\log\left(\frac{(2\pi)^{n}\pi^{n}}{\mathrm{vol}(X,\Delta)}\right)$
Interestingly, Lemma 3.2 reveals that the family of log Fano pairs $(\mathcal{X},\mathcal{D})$ appearing in the lemma may be explicitly expressed in terms of the algebro-geometric volume $\text{vol}(X,\Delta)$ in the same functional form as the one appearing in the previous upper and lower bounds:
$\frac{(\overline{-\mathcal{K}_{(\mathcal{X},\mathcal{D})}})^{n+1}}{(n+1)!}=\frac{1}{2}\text{vol}(X,\Delta)\log\left(\frac{be^{2a}}{\text{vol}(X,\Delta)}\right)$
with $a:=\frac{(\overline{-\mathcal{K}_{\mathcal{X}}})^{n+1}/(n+1)!}{\left(-K_{X}\right)^{n}/n!}$ and $b=\text{vol}(X).$
4. Hyperplane arrangements
In this Section we prove Theorem 1.3 concerning hyperplane arrangements. Recall that a log Fano pair $(X,\Delta)$ is called a log Fano hyperplane arrangement if $X=\mathbb{P}^{n}$ and $\Delta=\sum_{i=1}^{m}w_{i}H_{i}$ where $w_{i}\in\mathbb{Q}_{>0}$ and the $H_{i}$ are distinct hyperplanes. Furthermore we will call $(X,\Delta)$ simple normal crossing, abbreviated snc, if the support of $\Delta$ has simple normal crossings.
For an snc log Fano hyperplane arrangement, if $m=n+1$ and all the weights $w_{i}$ are equal, then $(X,\Delta)$ is a toric log-pair (see Lemma 3.2). The following lemma shows that for given hyperplanes $H_{1},...,H_{m}$ and a fixed volume $\text{vol}(X,\Delta)$, the “toric” weights form the vertices of the convex polytope of all weights $w_{i}$ corresponding to K-semistable $(X,\Delta).$
Lemma 4.1.
Fix $m\geq 1$ and a real number $0<D\leq(-K_{\mathbb{P}^{n}})^{n}=(n+1)^{n}.$ Let as before for a real $m$-tuple $w$, $\Delta=\sum_{i=1}^{m}w_{i}H_{i}$ for distinct hyperplanes $H_{i}$. Then the set of weights
$S=\{w\in\mathbb{R}^{n}:(-(K_{\mathbb{P}^{n}}+\Delta))^{n}=D\mathrm{\ and\ }(\mathbb{P}^{n},\Delta)\mathrm{\ is\ K-semistable}\}$
is either empty or $m\geq n+1$ and $S$ is a polytope with $\binom{n}{m}$ vertices given by any reordering of the tuple $w_{1}=w_{2}=...=w_{n+1}=\frac{1}{m}(n+1-D^{1/n})$ , $w_{l}=0$ $\forall l>n+1$.
Proof.
By [23], for $w\in\mathbb{R}^{n}$ and $\Delta=\sum_{i=1}^{m}w_{i}H_{i}$ , $(\mathbb{P}^{n},\Delta)$ is K-semistable and log Fano if and only if $w$ is in the convex set $C$ defined by the following inequalities:
00 0 $\leq w_{i}<1\ \forall i=1,...,m$
(4.1) $k$ $\sum_{i=1}^{m}w_{i}\geq(n+1)\sum_{j=1}^{k}w_{i_{j}}\ \forall 1\leq k\leq n\ \forall 1\leq i_{1}<...<i_{k}\leq m.$
Here it should be noted that in fact, it suffices to consider the second inequality for index combinations $i_{j}$ of length 1. The other inequalities for larger index combinations follows. Hence, K-semistability of $(\mathbb{P}^{n},\Delta)$ is equivalent to
00 0 $\leq w_{i}<1\ \forall i=1,...,m$
(4.2) $w_{i}$ $\leq\frac{1}{n+1}\sum_{j=1}^{m}w_{j}\ \forall i=1,...,m.$
Fix $m$ and $D$ as in the statement of the theorem. The goal is to understand the intersection of the above set with the set $\{w:-(K_{\mathbb{P}^{n}}+\Delta)=D\}.$ Note first that
$(-(K_{\mathbb{P}^{n}}+\sum_{i=1}^{m}w_{i}H_{i}))^{n}=(n+1-\sum_{i=1}^{m}w_{i})^{n}.$
Let $C:=n+1-D^{1/n}$, so that $\{w:-(K_{\mathbb{P}^{n}}+\Delta)=D\}=\{w:\sum_{i=1}^{m}w_{i}=C\}$. Thus with S defined as in 4.1
$S=\{w:$ $\sum_{i=1}^{m}w_{i}=C$
$0\leq w_{i}<1\ \forall i=1,...,m$
$w_{i}\leq\frac{C}{n+1}\ \forall i=1,...,m\}$
Observe that since $10\leq C<n+$ 1, the inequality $w_{i}<1$ is superfluous. After a convenient rescaling we get
$\frac{n+1}{C}S=\{w:$ $\sum_{i=1}^{m}w_{i}=n+1$
$0\leq w_{i}\leq 1\ \forall i=1,...,m\}.$
Clearly if $m<n+1$, $\frac{n+1}{C}S$ is empty. For $m\geq n+1$, any vertex of $\frac{n+1}{C}S$ is given by the intersection of $\frac{n+1}{C}S$ with some collection of the inequalities put to equality. But clearly all such points must be of the form of having $n+1$ ones and $m-(n+1)$ zeros. And on the other hand any such point is a vertex. ∎
Fixing the volume $\text{vol}(X,\Delta)$ is, when $X=\mathbb{P}^{n},$ tantamount to fixing the isomorphism class of the $\mathbb{Q}-$line bundle $-(K_{X}+\Delta)$ (since the rank of the Picard group of $\mathbb{P}^{n}$ is one). The following lemma shows that, in this case, the maximal height is convex with respect to the weights of $\Delta.$
Lemma 4.2.
Consider an arithmetic Fano variety $\mathcal{X}$ and a curve $t\mapsto(\mathcal{X},\mathcal{D}_{t})$ of arithmetic log Fano varieties where $\mathcal{D}_{t}=\sum_{i=1}^{m}w_{i}(t)\mathcal{D}_{i}$ for some $m\geq 1$, irreducible divisors $\mathcal{D}_{i}$ over $\mathbb{Z}$ and $w:[0,1]\rightarrow\mathbb{R}^{m}$ an affine function. Additionally assume that all the $\mathcal{D}_{t}$ are linearly equivalent, which equivalently means that $-(\mathcal{K}+\mathcal{D}_{t})$ isomorphic to $\mathcal{L}$ for a line bundle $\mathcal{L}\rightarrow\mathcal{X}$ independent of $t.$ Then the function $h:[0,1]\rightarrow]-\infty,\infty]$ defined as
[…]
Proof.
By assumption we can identify $-(\mathcal{K}+\mathcal{D}_{t})$ with $\mathcal{L}$ for a line bundle $\mathcal{L}$ independent of $\phi.$ Thus the height $h_{\text{$\phi$}}(\mathcal{X},\mathcal{D}_{t})$ for a fixed metric on $\mathcal{L}$ is independent of $t.$ Likewise, $h_{\text{can}}(\mathcal{X},\mathcal{D}_{t})$ coincides with $\hat{h}_{\text{can}}(\mathcal{X},\mathcal{D}_{t})$ up to multiplication by a constant independent of $t.$ Next, express
$\hat{h}_{\text{can}}(\mathcal{X},\mathcal{D}_{t})=\sup_{\phi}\hat{h}(\mathcal{X},\mathcal{D}_{t})+\frac{1}{2}\log\int_{X}\mu_{(\phi,\mathcal{D}_{t})},$
where the sup ranges over all continuous psh metrics on $\mathcal{L}.$ Introducing an arbitrary volume form $\mathrm{d}V$ on $X$ we can rewrite
[…]
By Hölder’s inequality this expression is convex in $t,$ since $w_{i}(t)$ is affine. It is even strictly convex since the $D_{i}$ are distinct. This means that $\hat{h}_{\text{can}}(\mathcal{X},\mathcal{D}_{t})$ is the supremum over a set independent of $t$, of a collection of strictly convex functions and thus is itself, strictly convex. ∎
4.1. Conclusion of the proof of Theorem 1.3
In the following we will use the notation $\mathcal{D}_{w}=\sum_{i=1}^{m}w_{i}H_{i}$ for $w\in\mathbb{R}^{m}$ for fixed hyperplanes $H_{i}$ defined over $\mathbb{Z}$. Let $(\mathbb{P}_{\mathbb{Z}}^{n},\mathcal{D}_{w^{\prime}})$ be a K-semistable snc log Fano hyperplane arrangement. Define for brevity $d:=(-(K_{\mathbb{P}^{n}}+\Delta(w^{\prime})))^{n}$. Set, as in Lemma 4.1, $S=\{w\in\mathbb{R}^{n}:(-(K_{\mathbb{P}^{n}}+\Delta_{w}))^{n}=d\mathrm{\ and\ }(\mathbb{P}_{\mathbb{Z}}^{n},\mathcal{D}_{w})\mathrm{\ is\ K-semistable}\}$. Consider the function $h(w)$ defined by
$h(w)=h_{\text{can}}(\mathbb{P}_{\mathbb{Z}}^{n},\mathcal{D}_{w})$
Restricted to the convex set $S$, $h|_{S}$ is convex by Lemma 4.2. Next by Lemma 4.1, $S$ is the convex hull of weight vectors $(w^{k})_{k=1,...,\binom{m}{n+1}}$, each corresponding to toric log Fano pairs, all equivalent to $(\mathbb{P}_{\mathbb{Z}}^{n},(1-t)\mathcal{D}_{0})$. Here $\mathcal{D}_{0}$ is the toric standard anti-canonical divisor and $t$ is the unique number such that $(-(K_{\mathbb{P}^{n}}+(1-t)\mathcal{D}_{0})^{n}=d$. By Jensens inequality it follows directly from expressing $w$ as a convex combination of the $w^{k}$, i.e. $w=\sum_{k=1}^{\binom{m}{n+1}}\lambda_{k}w^{k}$, that
$h(w)\leq\sum_{k=1}^{\binom{m}{n+1}}\lambda_{k}h(w^{k})=h(w^{1})=h_{\text{can}}(\mathbb{P}_{\mathbb{Z}}^{n},(1-t)\mathcal{D}_{0})$
with equality iff $\Delta_{w}$ is toric. We have thus reduced to a toric case, which we have already handled. Specifically, the bound 1.4 follows directly from Lemma 3.2. For Theorem 1.3, recall that it was observed in the proof of Lemma 3.3 that the volume dependent bound in 1.4 is increasing with volume, so that a universal bound is given for maximal volume, i.e. when $\Delta=0$, yielding the result.
5. Diagonal hypersurfaces
In this section we will deduce Theorem 1.4 from the results in the previous sections. The starting point of the proof is the following analytic representation of the height:
Lemma 5.1.
(Restriction formula) Let $\mathcal{X}$ be the subscheme of $\mathbb{P}_{\mathbb{Z}}^{n+1}$ cut out by a homogeneous polynomial $s$ of degree $d$ with integer coefficients and $\phi$ a continuous psh metric on $\mathcal{O}(d)\rightarrow\mathbb{P}_{\mathbb{C}}^{n+1}.$ Then the height $h_{\phi}(\mathcal{X}_{d},\mathcal{O}(d))$ of the restriction of $(\mathcal{O}(d),\phi)$ to $\mathcal{X}$ may be expressed as
$\frac{2h_{\phi}(\mathcal{X}_{d},\mathcal{O}(d))}{(n+1)!}=(n+2)\mathcal{E}(\phi,d\phi_{0})+\int_{\mathbb{P}^{n+1}}\log\left(\left\|s\right\|_{\phi}^{2}\right)\frac{(dd^{c}\phi)^{n+1}}{(n+1)!},$
where $\phi_{0}$ is the Weil metric on $\mathcal{O}(1)$ and $\mathcal{E}$ is the functional defined by formula 2.7, corresponding to $\mathcal{O}(d)\rightarrow\mathbb{P}^{n+1}.$
Proof.
This is well-known, but for completeness we provide a proof. Consider first the general situation where $\mathcal{X}$ is a subscheme (of relative dimension $n)$ of a regular projective flat scheme $\mathcal{Y}$ cut out by a section $s$ of a relatively ample line bundle $\mathcal{L}\rightarrow\mathcal{Y}.$ Then, given a metric $\phi$ on the complexification $L$ of $\mathcal{L}\rightarrow\mathcal{Y},$ the restriction formula for arithmetic intersection numbers [7, Prop 2.3.1] gives
(5.1) $(\mathcal{L},\phi)^{n+2}\cdot\mathcal{Y}=(\mathcal{L},\phi)^{n+1}\cdot\mathcal{X}-\int_{Y}\log||s||_{\phi}(dd^{c}\phi)^{n+1}$
In particular, setting $\mathcal{Y}=\mathbb{P}_{\mathbb{Z}}^{n+1}$ and $\mathcal{L}=\mathcal{O}(d)$ gives and
$\frac{2h_{\phi}(\mathcal{X}_{d},\mathcal{O}(d))}{(n+1)!}=(n+2)\frac{h_{\phi}(\mathbb{P}_{\mathbb{Z}}^{n+1},\mathcal{O}(d))}{(n+2)!}+\int\log||s||_{\phi}^{2}\frac{(dd^{c}\phi)^{n+1}}{(n+1)!}.$
The proof is thus concluded by invoking the well-known fact that $h_{\phi}(\mathbb{P}_{\mathbb{Z}}^{n+1},\mathcal{O}(d))/(n+2)!=\mathcal{E}_{\mathbb{P}^{n+1}}(\phi,\phi_{0}).$ For example, this is a special case of the toric formula in [1, formula 3.7]. ∎
In general, if $\mathcal{X}$ is subscheme of $\mathbb{P}_{\mathbb{Z}}^{n+1}$ of codimension one, then $\mathcal{K}_{\mathcal{X}}$ is well-defined as line bundle over $\mathcal{X}.$ More precisely, by the adjunction formula, there is an isomorphism of line bundles over $\mathbb{Z},$
$\mathcal{K}_{\mathcal{X}}\simeq\left(\mathcal{K}_{\mathbb{P}_{\mathbb{Z}}^{n+1}}-\mathcal{O}(\mathcal{I}/\mathcal{I}^{2})\right)_{\mathcal{X}},$
where $\mathcal{I}$ is the ideal sheaf cutting out $\mathcal{X}$ [27, formula 1.6.2, page 8]. In particular, if $\mathcal{X}$ is cut out by a homogeneous polynomial $s$ of degree $d,$ then
(5.2) $-\mathcal{K}_{\mathcal{X}}\simeq-\mathcal{K}_{\mathbb{P}_{\mathbb{Z}}^{n+1}}-\mathcal{O}(d)\simeq\mathcal{O}(n+2-d).$
Hence, $-\mathcal{K}_{\mathcal{X}}$ is relatively ample iff $d\leq n+1.$ Now assume that the complex variety $X$ defined by the complex points of $\mathcal{X}$ is non-singular. Then, by the adjunction isomorphism 5.2, a metric $\phi_{0}$ on $\mathcal{O}(n+2-d)|_{X}$ may be identified with a metric on $-K_{X}.$
5.1. Reduction to Fermat hypersurfaces
Given integers $a_{i}$ consider the subscheme $\mathcal{X}_{a}$ of $\mathbb{P}_{\mathbb{Z}}^{n+1}$ cut out by the homogeneous polynomial
$s_{a}:=\sum_{i=0}^{n+1}a_{i}x_{i}^{d}.$
Denote by $X_{a}$ the corresponding complex variety, which is non-singular and consider the map
(5.3) $F_{a}(x):=(a_{0}^{-1/d}x_{0},...,a_{n+1}^{-1/d}x_{n+1}):\,\,\mathbb{C}^{n+2}\rightarrow\mathbb{C}^{n+2}.$
We will identify the map $F_{a}$ with an automorphism of $\mathbb{P}^{n+1},$ admitting a standard lift to the total space of $\mathcal{O}(1)\rightarrow\mathbb{P}^{n+1}.$ We can then express
$X_{a}=F_{a}(X_{1}).$
Proposition 5.2.
Let $k$ be a positive integer and $\phi$ a metric on $\mathcal{O}(n+2-d)|_{X_{a}}.$ Then
$2\hat{h}_{\phi}(\mathcal{X}_{a},\mathcal{O}(n+2-d))+\log\int_{X_{a}}e^{-\phi}=$
$=2\hat{h}_{F_{a}^{*}\phi}(\mathcal{X}_{1},\mathcal{O}(n+2-d))+\log\int_{X_{1}}e^{-F_{a}^{*}\phi}+(\frac{n+2-d}{(n+1)}-1)d^{-1}\sum_{i}\log(|a_{i}|^{2}).$
The proof of the proposition follows from combining the following two lemmas:
Lemma 5.3.
Let $k$ be a positive integer and $\phi$ a metric on $\mathcal{O}(k).$ Then
$2\widehat{h}_{\phi}(\mathcal{X}_{a},\mathcal{O}(k))=2\widehat{h}_{F_{a}^{*}\phi}(\mathcal{X}_{1},\mathcal{O}(k))+\frac{1}{d}\sum_{i}\log(|a_{i}|^{2})\frac{k}{(n+1)}$
Proof.
We will use that any continuous psh metric $\phi$ on $\mathcal{O}(k)|_{X}$ is the restriction of a continuous psh metric on $\mathcal{O}(k)\rightarrow\mathbb{P}^{n+1},$ that we shall denote by the same symbol $\phi$ [14]. First consider the case when $k=d$ and denote by $s_{a}$ the section of $\mathcal{O}(d)$ cutting out the scheme $\mathcal{X}_{a}.$ By the restriction formula (Lemma 5.1)
$\frac{2}{(n+1)!}h_{\phi}(\mathcal{X}_{a},\mathcal{O}(k))=(n+2)\mathcal{E}_{\mathbb{P}^{n+1}}(\phi,d\phi_{0})+\int_{\mathbb{P}^{n+1}}\log|s_{a}|_{\phi}^{2}MA(\phi).$
Rewriting
$\int_{\mathbb{P}^{n+1}}\log|s_{a}|_{\phi}^{2}MA(\phi)=\int_{\mathbb{P}^{n+1}}\log|(F_{a}^{-1})^{*}s_{1}|_{\phi}^{2}MA(\phi)=\int_{\mathbb{P}^{n+1}}\log|s_{1}|_{F_{a}^{*}\phi}^{2}MA(F_{a}^{*}\phi),$
thus reveals that
$\frac{2}{(n+1)!}h_{F_{a}^{*}\phi}(X_{1},\mathcal{O}(k))-\frac{2}{(n+1)!}h_{\phi}(X_{a},\mathcal{O}(k))=(n+2)\left(\mathcal{E}_{\mathbb{P}^{n+1}}(F_{a}^{*}\phi,d\phi_{0})-\mathcal{E}_{\mathbb{P}^{n+1}}(\phi,d\phi_{0})\right)$
Now, denote by $G_{a}$ the standard lift of $F_{a}$ from $X$ to $-K_{X}$ and its tensor powers. We can then express
$G_{a}^{*}\phi=F_{a}^{*}\phi+c_{a},\,\,\,c_{a}:=\frac{k}{(n+2)}\frac{1}{d}\sum_{i}\log(|a_{i}|^{2})$
Indeed,
(5.4) $G_{a}^{*}(e^{-\frac{n+2}{k}\phi}dzd\bar{z})=(e^{-\frac{n+2}{k}F_{a}^{*}\phi})F_{a}^{*}(dzd\bar{z})=(e^{-\frac{n+2}{k}F_{a}^{*}\phi})\prod_{i}|a_{i}|^{-2/d}(dzd\bar{z}).$
Hence,
$2h_{F_{a}^{*}\phi}(\mathbb{P}^{n+1},\mathcal{O}(k))=2h_{G_{a}^{*}\phi-c_{a}}(\mathbb{P}^{n+1},\mathcal{O}(k))=2h_{G_{a}^{*}\phi}(\mathbb{P}^{n+1},\mathcal{O}(k))-c_{a}\frac{k^{n+1}}{(n+1)!}$
But, by Lemma 7.1 in the appendix
$\mathcal{E}_{\mathbb{P}^{n+1}}(G_{a}^{*}\phi,d\phi_{0})=\mathcal{E}_{\mathbb{P}^{n+1}}(\phi,d\phi_{0})$
Hence, $\frac{2}{(n+1)!}h_{F_{a}^{*}\phi}(X_{1},\mathcal{O}(k))-\frac{2}{(n+1)!}h_{\phi}(X_{a},\mathcal{O}(k))=$
$=(n+2)\left(\mathcal{E}_{\mathbb{P}^{n+1}}(F_{a}^{*}\phi,d\phi_{0})-\mathcal{E}_{\mathbb{P}^{n+1}}(\phi,d\phi_{0})\right)=-(n+2)c_{a}\frac{k^{(n+1)}}{(n+1)!}$
As a consequence,
$2\hat{h}_{F_{a}^{*}\phi}(X_{1},\mathcal{O}(k))-2\hat{h}_{\phi}(X_{a},\mathcal{O}(k))=-(n+2)c_{a}\frac{k^{n+1}}{(n+1)!}.\frac{1}{dk^{n}/n!}=-(n+2)c_{a}\frac{k}{d(n+1)}=$
$=-\frac{1}{d}\sum_{i}\log(|a_{i}|^{2})(n+2)\frac{k}{(n+2)}\frac{k}{d(n+1)}=-\frac{1}{d}\sum_{i}\log(|a_{i}|^{2})\frac{k^{2}}{d(n+1)}$
Since we have assumed that $k=d$ this means that
$2\widehat{h}_{\phi}(X_{a},\mathcal{O}(k))-2\widehat{h}_{F_{a}^{*}\phi}(X_{1},\mathcal{O}(k))=\frac{1}{d}\sum_{i}\log(|a_{i}|^{2})\frac{d}{(n+1)}$
Finally, for any given integer $k$ we can express $k=d\lambda$ for $\lambda=k/d$ and use the basic scaling property
$\widehat{h}_{\lambda\phi}(\mathcal{X},\lambda\mathcal{L})=\lambda\widehat{h}_{\phi}(\mathcal{X},\mathcal{L}),$
to get
$2\widehat{h}_{\phi}(X_{a},\mathcal{O}(k))-2\widehat{h}_{F_{a}^{*}\phi}(X_{1},\mathcal{O}(k))=\frac{1}{d}\sum_{i}\log(|a_{i}|^{2})\frac{k}{(n+1)}$
which concludes the proof. ∎
Lemma 5.4.
Given a metric $\phi$ on $\mathcal{O}(n+2-d)$ we have that
$\log\int_{X_{a}}e^{-\phi}=\log\int_{X_{1}}e^{-F_{a}^{*}\phi}-d^{-1}\sum_{i}\log(|a_{i}|^{2}).$
Proof.
It will be convenient to lift the integrals to the corresponding affine cones. To this end let $X$ be the non-singular hypersurface of $\mathbb{P}^{n+1},$ cut out by a given homogeneous polynomial $s$ in $\mathbb{C}^{n+2}.$ Let $A_{X}$ be the affine cone over $X:$
$A_{X}=\{s=0\}\subset\mathbb{C}^{n+2}$
Let $\Omega_{s}$ be the holomorphic top form on $A_{X}-\{0\}$ defined by the relation
$\Omega_{s}\wedge ds=dz\,\,\,\text{on $Y$},\,\,\,\,dz:=dz_{0}\wedge\cdots\wedge dz_{n+1}.$
Denote by $\delta$ interior multiplication with the generator of the standard $\mathbb{R}_{+}-$action, acting by scalings, on $\mathbb{C}^{n+2}.$ Assume now that $X$ is Fano. A given metric $\phi$ on $-K_{X}$ then induces a one-homogeneous function $r$ on $A$ (using the adjunction isomorphism 5.2 and by identifying $\mathbb{C}^{n+2}-\{0\}$ with the complement of the zero-section in $\mathcal{O}(1)^{*}\rightarrow X).$ Moreover, lifting the adjunction isomorphism 5.2 to $A_{X}$ yields the following well-known formula (which applies in the general setup of Fano varieties over local fields; cf. [38, Lemma 4.2.2]):
$\int_{X}e^{-\phi}=c\int_{\{r=1\}\cap A_{X}}\delta(\Omega_{s}\wedge\overline{\Omega_{s}})$
for a non-zero constant $c$ only depending on $n$ and $d.$ Hence, recalling that $F$ denotes the scaling map in formula 5.3 on $\mathbb{C}^{n+2},$
$\int_{X}e^{-\phi}=c\int_{\{F^{*}r=1\}\cap F^{*}A_{X}}F^{*}(\delta(\Omega_{s}\wedge\overline{\Omega_{s}}).$
Using that the $\mathbb{R}_{+}-$action commutes with $F$ thus gives
$\int_{X}e^{-\phi}=c\int_{\{F^{*}r=1\}\cap A_{F^{*}X}}\delta(F^{*}(\Omega_{s})\wedge\overline{F^{*}\Omega_{s}}).$
Applying $F^{*}$ to the defining relation for $\Omega_{s}$ yields
$F^{*}\Omega_{s}\wedge d(F^{*}f)=F^{*}dz\,\,\,\text{on $Y$}.$
Since $F^{*}dz=a_{0}^{-1/d}\cdots a_{n+1}^{-1/d}dz$ this shows that $F^{*}\Omega_{s}=\Omega_{F^{*}s}a_{0}^{-1/d}\cdots a_{n+1}^{-1/d},$ which concludes the proof. ∎
It follows directly from Proposition 5.2 that
(5.5) $h_{\text{can }}(\mathcal{X}_{a})=h_{\text{can }}(\mathcal{X}_{1})+(n+1)(n+2-d)^{n}d\left(\frac{n+2-d}{(n+1)}-1\right)d^{-1}\sum_{i}\log(|a_{i}|^{2}).$
In particular, since $n+2-d\leq n+1,$ this means that $h_{\text{can }}(\mathcal{X}_{a})\leq h_{\text{can }}(\mathcal{X}_{1})$ and thus the proof of Theorem 1.4 is reduced to the case of the Fermat hypersurface $\mathcal{X}_{1}.$
Remark 5.5.
More generally, let $\mathcal{X}$ be a hypersurface in $\mathbb{P}_{\mathbb{Z}}^{n+1}$ cut-out by a homogeneous polynomial $s$ of degree $d$ of the form $T^{*}s_{1}$ where $T\in GL(n+2,\mathbb{C}).$ Then formula 5.5 can be generalized as follows (as shown in essentially the same manner as before):
$h_{\text{can }}(\mathcal{X})=h_{\text{can }}(\mathcal{X}_{1})+(n+1)(n+2-d)^{n}d\left(\frac{n+2-d}{(n+1)}-1\right)\log(|\det T|^{2}).$
5.2. Reduction to log hyperplane arrangements
Fix a degree $d(\leq n+1)$ and denote by $\mathcal{X}$ the corresponding Fermat hypersurface. The Fermat hypersurface of degree one will be denoted by $\mathcal{Y}.$ We will next express the canonical height $h_{\text{can }}(\mathcal{X})$ in terms of the canonical height $h_{\text{can }}(\mathcal{Y},\mathcal{D})$ where $\mathcal{D}$ is the divisor on $\mathcal{Y}$ defined by
$\mathcal{D}=(1-1/d)\left[x_{0}=0\right]+...+(1-1/d)\left[x_{n+1}=0\right],$
where $x_{i}$ denotes the homogenous coordinates on $\mathbb{P}_{\mathbb{Z}}^{n+1}$ restricted to $\mathcal{Y}.$
Proposition 5.6.
Denote by $\mathcal{X}$ the Fermat hypersurface of a given degree $m(\leq n+1)$ and by $\mathcal{Y}$ the Fermat hypersurface of degree one, endowed with the divisor $\mathcal{D}.$ Then
(5.6) $\hat{h}_{\text{can }}(\mathcal{X})=\hat{h}(\mathcal{Y},\mathcal{D})-\frac{1}{2}\log\frac{V(X)}{V(Y,\Delta)}$
Proof.
By the adjunction formula we have isomorphisms $-K_{\mathcal{X}}\simeq(n+2-m)\mathcal{O}(1)|_{\mathcal{X}}$ and $-\mathcal{K}_{(\mathcal{Y},\Delta)}\simeq(n+2-m)(1/m)\mathcal{O}(1)|_{\mathcal{Y}}.$ Consider the following morphism:
$F:\,\mathbb{C}^{n+2}\rightarrow\mathbb{C}^{n\text{+2}},\,\,\,(x_{0},...,x_{n+1})\mapsto(x_{0}^{m},...,x_{n+1}^{m}),$
which induces a map $\mathbb{P}^{n+1}\rightarrow\mathbb{P}^{n\text{+1}}$ and a lift to $\mathcal{O}(1),$ which is naturally defined over $\mathbb{Z}$ and satisfies $F^{*}\mathcal{O}(1)\simeq m\mathcal{O}(1).$ In particular, it induces a morphism
$F:\,\mathcal{X}\rightarrow\mathcal{Y},\,\,\,\,F^{*}(-\mathcal{K}_{(\mathcal{Y},\Delta)})\simeq-K_{\mathcal{X}}$
under the adjunction isomorphisms. We recall that, by basic functorial properties of heights,
(5.7) $\hat{h}(\mathcal{X},F^{*}\mathcal{L})=\hat{h}(\mathcal{Y},\mathcal{L}).$
In fact, in this case this formula follows directly from the analytic representation of the height in Lemma 5.1, using that $F$ preserves the Weil metric $\phi_{0}.$ In particular, setting $\mathcal{L}:=(n+2-m)(1/m)\mathcal{O}(1)_{|\mathcal{Y}}$ and using the adjunction isomorphisms yields $\hat{h}(\mathcal{X},-K_{\mathcal{X}},F^{*}\phi)=\hat{h}(\mathcal{Y},-\mathcal{K}_{(\mathcal{Y},\Delta)},\phi).$ Thus, all that remains is to show that
$\int_{X}\mu_{F^{*}\phi}=m^{-(n+1)}\int_{Y}\mu_{\phi},$
where we have used that $F^{*}\phi$ induces a metric on $-K_{X}$ and $\phi$ induces a metric on $-K_{(Y,\Delta)}.$ Since $F$ has topological degree $m^{(n+1)}$ we have $F_{*}[X_{m}]=m^{(n+1)}[Y]$ as homology classes and thus it will be enough to show that
$F^{*}\mu_{\phi}=m^{2(n+1)}\mu_{\phi}.$
To this end consider the affine piece $\mathbb{C}^{n+1}$ of $\mathbb{P}^{n+1}$ where $x_{0}\neq 0.$ Setting $z_{i}=x_{i}/x_{0}$ for $i=1,...,n+1)$ we can, locally, parametrize $X$ by the coordinates $z_{1},...,z_{n}.$ In these coordinates a metric $\psi$ on the restriction of $\mathcal{O}(1)$ to $X$ induces, by the adjunction isomorphism, a metric on $-K_{X}$ and thus a measure on $X$ locally expressed as
(5.8) $\mu_{\psi}=\frac{e^{-(n+2-m)\psi}}{\left(m|z_{n+1}|^{m-1}\right)^{2}}\frac{i}{2}\mathrm{d}z_{1}\wedge\mathrm{d}\bar{z}_{1}\cdots\frac{i}{2}\mathrm{d}z_{n}\wedge\mathrm{d}\bar{z}_{n}.$
To see this, recall that, by definition,
$\mu_{\psi}:=\left\|\mathrm{d}z_{1}\wedge\cdots\wedge\mathrm{d}z_{n}\right\|_{\psi}^{-2}\frac{i}{2}\mathrm{d}z_{1}\wedge\mathrm{d}\bar{z}_{1}\cdots\frac{i}{2}\mathrm{d}z_{n}\wedge\mathrm{d}\bar{z}_{n}.$
In the affine piece $\mathbb{C}^{n+1}$ of $\mathbb{P}^{n+1}$ we can express
$s=fx_{0}^{\otimes 2},\,\,\,\,f=1+\sum_{i=1}^{n}z_{i}^{m}+z_{n+1}^{m}$
By the adjunction isomorphism 5.2 we have
$\left\|\mathrm{d}z_{1}\wedge\cdots\wedge\mathrm{d}z_{n}\right\|:=\left\|\mathrm{d}z_{1}\wedge\cdots\wedge\mathrm{d}z_{n}\wedge\mathrm{d}s\right\|=\left\|\mathrm{d}z_{1}\wedge\cdots\wedge\mathrm{d}z_{n}\wedge\mathrm{d}f\right\|\left\|x_{0}^{\otimes m}\right\|.$
Since $\mathrm{d}z_{1}\wedge\cdots\wedge\mathrm{d}z_{n}\wedge\mathrm{d}f=\mathrm{d}z_{1}\wedge\cdots\wedge\mathrm{d}z_{n}\wedge\mathrm{d}z_{n+1}\partial f/\partial z_{n+1}$ this means that
$\left\|\mathrm{d}z_{1}\wedge\cdots\wedge\mathrm{d}z_{n}\right\|^{2}:=\left|\frac{\partial f}{\partial z_{n+1}}\right|^{2}\left\|\mathrm{d}z_{1}\wedge\cdots\wedge\mathrm{d}z_{n}\wedge\mathrm{d}z_{n+1}\right\|\left\|x_{0}^{\otimes m}\right\|=e^{(n+2)\psi}e^{-m\psi},$
giving
(5.9) $\mu_{\phi}=\frac{e^{-(n+2-m)\psi}}{|\partial f/\partial z_{n+1}|^{2}}\frac{i}{2}\mathrm{d}z_{1}\wedge\mathrm{d}\bar{z}_{1}\cdots\frac{i}{2}\mathrm{d}z_{n}\wedge\mathrm{d}\bar{z}_{n},$
which proves 5.8. Likewise, we can parametrize the affine piece of $Y$ by the coordinates $z_{1},...,z_{n}.$ A given metric $\phi$ on the restriction of $\mathcal{O}(1)$ to $Y$ induces, by the adjunction isomorphism a measure (defined with respect to the divisor $\mathcal{D})$
$\mu_{\phi}=e^{-(n+2-m)m^{-1}\phi}\left|z_{n+1}\right|^{-2(1-1/m)}\left|z_{1}\right|^{-2(1-1/m)}\cdots\left|z_{n}\right|^{-2(1-1/m)}\frac{i}{2}\mathrm{d}z_{1}\wedge\mathrm{d}\bar{z}_{1}\cdots\frac{i}{2}\mathrm{d}z_{n}\wedge\mathrm{d}\bar{z}_{n}.$
Since $F^{*}z_{i}=z_{i}^{m}$ this means that
$F^{*}\mu_{\phi}=e^{-(n+2-m)f^{*}(m^{-1}\phi)}\left|z_{n+1}\right|^{-2(m-1)}\left|z_{1}\right|^{-2(m-1)}\cdots\left|z_{n}\right|^{-2(m-1)}\frac{i}{2}\mathrm{d}(z_{1}^{m})\wedge\mathrm{d}(\bar{z}_{1}^{m})\cdots\frac{i}{2}\mathrm{d}(z_{n}^{m})\wedge\mathrm{d}(\bar{z}_{n}^{m}).$
Finally, since $\mathrm{d}(z^{m})=mz^{m-1}$ this proves the desired identity 5.6, using the representation 5.9 with $\psi=F^{*}(m^{-1}\phi).$
5.3. Conclusion of the proof of Theorem 1.4
The affine projection $(x_{0},..,x_{n+1})\mapsto(x_{0},...,x_{n})$ induces an isomorphism from $\mathcal{Y}$ to $\mathbb{P}_{\mathbb{Z}}^{n},$ identifying $(\mathcal{Y},\mathcal{D})$ with a hyperplane arrangement $(\mathbb{P}_{\mathbb{Z}}^{n},\mathcal{D})$ with simple normal crossings. It follows readily from the definition of $\mathcal{D}$ and the criterion 4.1 that $(\mathbb{P}_{\mathbb{Z}}^{n},\mathcal{D})$ is K-semistable. Hence, combining Proposition 5.6 with refined bound following the statement of Theorem 1.3 yields
$\hat{h}_{\text{can }}(\mathcal{X})\leq\hat{h}_{\text{can }}(\mathbb{P}_{\mathbb{Z}}^{n},\mathcal{D}_{t})-\frac{1}{2}\log\frac{V(X)}{V(\mathbb{P}^{n},\Delta_{t})},$
where $\mathcal{D}_{t}$ is the toric divisor on $\mathbb{P}_{\mathbb{Z}}^{n}$ such that $(\mathbb{P}_{\mathbb{Z}}^{n},\mathcal{D}_{t})$ is K-semistable and $V(\mathbb{P}^{n},\Delta_{t})=V(\mathbb{P}^{n},\Delta).$ The explicit formula for $\hat{h}_{\text{can }}(\mathbb{P}_{\mathbb{Z}}^{n},\mathcal{D}_{t})$ thus yields
$\hat{h}_{\text{can }}(\mathcal{X})\leq\hat{h}_{\text{can }}(\mathbb{P}_{\mathbb{Z}}^{n})-\frac{1}{2}\log\frac{V(X)}{V(\mathbb{P}^{n})}.$
Multiplying both sides with $V(X)$ reveals that
$h_{\text{can }}(\mathcal{X})\leq\lambda h_{\text{can }}(\mathbb{P}_{\mathbb{Z}}^{n})-\frac{1}{2}V(X)\log\lambda,\,\,\,\lambda:=V(X)/V(\mathbb{P}^{n}).$
Since $\lambda\in]0,1[$ it thus follows from Lemma 3.2 that the right hand side above is increasing with respect to $\lambda$ and thus maximal when $\lambda=1,$ giving $h_{\text{can }}(\mathcal{X},\mathcal{D})\leq h_{\text{can }}(\mathbb{P}_{\mathbb{Z}}^{n}).$ Moreover, the equality is strict if $d\geq 2$ since then $\lambda<1.$
6. Arithmetic log surfaces
In this section we will prove Theorem 1.5. As explained in Section 1.1.4, by Theorem 1.3, the proof is reduced to showing that for any fixed metric on $-K_{(X,\Delta)}:$
  • The canonical integral model $(\mathcal{X}_{c},\mathcal{D}_{c};-\mathcal{K}_{(\mathcal{X}_{c},\mathcal{D}_{c})})$ of $(X,\Delta;-K_{(X,\Delta)})$ obtained by setting $\mathcal{X}_{c}=\mathbb{P}_{\mathbb{Z}}^{1}$ and taking $\mathcal{D}_{c}$ to be the Zariski closure of $\Delta_{\mathbb{Q}}$ in $\mathcal{X}_{c}$ minimizes $\mathcal{M}_{(\mathcal{X},\mathcal{D})}(\overline{\mathcal{L}})$ over all integral models $(\mathcal{X},\mathcal{D};\mathcal{L})$ of $(X,\Delta;-K_{(X,\Delta)}$)
  • When $\mathcal{D}=0$ the minimum is uniquely attained for $\mathcal{X}=\mathbb{P}_{\mathbb{Z}}^{1},$ up to isomorphisms over $\mathbb{Z}.$
6.1. Preliminaries on log canonical thresholds
Following [27, 46] a log pair $(\mathcal{X},\mathcal{D})$ is said to be log canonical (lc) if for any normal blow-up morphism $p:\mathcal{Y}\rightarrow\mathcal{X}$
$\mathcal{K}_{\mathcal{Y}/\mathcal{X}}-p^{*}\mathcal{D}=\sum_{i}a_{i}E_{i},\,\,\,a_{i}\geq-1,\,\,\,\mathcal{K}_{\mathcal{Y}/\mathcal{X}}:=\mathcal{K}_{\mathcal{Y}}-p^{*}\mathcal{K}_{\mathcal{X}}$
where the prime divisor $E_{i}$ is either an exceptional divisor of $p$ or the proper transform of a component of $\mathcal{D}.$ The log canonical threshold of a $\mathbb{Q}-$divisor $F$ on $\mathcal{X}$ with respect to the log pair $(\mathcal{X},\mathcal{D})$ is defined by
lct $(\mathcal{X},\mathcal{D};F):=\sup_{t>0}\left\{t:\,(\mathcal{X},tF+\mathcal{D})\,\text{is lc}\right\}$
The following lemma follows readily from the definition:
Lemma 6.1.
For any normal blow-up morphism $p:\mathcal{Y}\rightarrow\mathcal{X}$
lct $(\mathcal{X},\mathcal{D};F)\leq\inf_{i}\frac{a_{i}-b_{i}+1}{c_{i}}$
where $a_{i},b_{i}$ and $c_{i}$ denote the order of vanishing along the $p-$exceptional prime divisor $E_{i}$ of $\mathcal{K}_{\mathcal{Y}/\mathcal{X}},$ $p^{*}\mathcal{D}$ and $p^{*}F,$ respectively and $i$ ranges over all $p-$exceptional prime divisors.
6.2. Preparations for the proof of Theorem 1.5
The following result is a logarithmic generalization of [37, Thm 2.14 (3)] (in the case of arithmetic surfaces).
Lemma 6.2.
Let $(X,D)$ be a log Fano curve over $\mathbb{C}$ and $(\mathcal{X},\mathcal{D})$ an arithmetic log Fano model for $(X,D)$ such the fibers $\mathcal{X}_{b}$ of $\mathcal{X}$ are reduced and irreducible and the divisor $\mathcal{D}$ is horizontal (i.e. $\mathcal{D}$ is the Zariski closure of $D)$. Assume that
$\alpha(\mathcal{X},\mathcal{D}):=\inf_{b,F}\text{lct $(\mathcal{X},\mathcal{D}+\mathcal{X}_{b};\,F)$}\geq 1/2$
where the inf runs over all effective $\mathbb{Q}-$divisors $F$ on $\mathcal{X}$ linearly equivalent to $-\mathcal{K}_{(\mathcal{X},\mathcal{D})}$ and closed points $b$ in the base $\mathcal{B}:=\text{Spec $\mathbb{Z}$ }$ such that $F$ does not contain the support of $\mathcal{X}_{b}.$ Then
$\frac{1}{2}\overline{\mathcal{L}}^{2}+\overline{\mathcal{K}}_{(\mathcal{X}^{\prime},\mathcal{D})}\cdot\overline{\mathcal{L}}\geq\frac{1}{2}\overline{\mathcal{L}}^{2}+\overline{\mathcal{K}}_{(\mathcal{X},\mathcal{D})}\cdot\overline{\mathcal{L}}$
for any relatively ample model $(\mathcal{X}^{\prime},\mathcal{D}^{\prime};\mathcal{L}^{\prime})$ of $(X,D;-K_{(X,D)})$ and given metrics on $-K_{(X,D)}$ and $L.$
In the proof we fix once and for all metrics on $-K_{(X,D)}$ and $L$ and set
(6.1) $\mathcal{M}_{(\mathcal{X},\mathcal{D})}(\mathcal{L}):=\frac{1}{2}\overline{\mathcal{L}}^{2}+\overline{\mathcal{K}}_{(\mathcal{X},\mathcal{D})}\cdot\overline{\mathcal{L}}$
for the corresponding metrized lines bundles $\overline{\mathcal{L}}$ and $\overline{\mathcal{K}}_{(\mathcal{X},\mathcal{D})}.$ Thus $\mathcal{M}_{(\mathcal{X},\mathcal{D})}(\mathcal{L})$ specializes to the arithmetic log Mabuchi functional $\mathcal{M}_{(\mathcal{X},\mathcal{D})}(\overline{\mathcal{L}})$ (formula 1.5) precisely when the metric on $\overline{\mathcal{K}}_{(\mathcal{X},\mathcal{D})}$ is the one induced from the curvature form $\omega$ of $\overline{\mathcal{L}}.$ But here it will be convenient to consider the present more general setup.
Proof when $D$ is trivial
To fix ideas we first consider the case when $D$ is trivial. Set $\mathcal{B}:=\text{Spec $\mathbb{Z}$ }$ and $\mathcal{L}:=-\mathcal{K}_{\mathcal{X}}.$ To simplify the notation we will remove the bar indicating the metric in the notation for the arithmetic intersection numbers. Anyhow, all the arithmetic intersections will be computed over the closed points $b$ in the base $\mathcal{B}$ and are thus independent of the choice of metric (since they are proportional to the algebraic intersections on the scheme $\pi^{-1}(b)$ over the residue field of $b$). If $F_{1}$ and $F_{2}$ are $\mathbb{Q}-$divisors we will write $F_{1}\geq F_{2}$ if $F_{1}-F_{2}$ is effective.
Step 1: It is enough to consider the case of a relatively semi-ample model of the form $(\mathcal{X}^{\prime},\mathcal{L}^{\prime})=(\mathcal{Y},p^{*}\mathcal{L}-E)$ where $p:\mathcal{Y}\rightarrow\mathcal{X}$ is the blow-up along a closed subscheme $\mathcal{Z}$ of $\mathcal{X}$ and $E$ is an effective $p-$exceptional divisor on $\mathcal{Y}$ whose support contains all the $p-$exceptional prime divisors and such that for any $b\in\pi(\mathcal{Z})$ $p^{*}\mathcal{L}-E$ admits a global section $s_{b}$ not vanishing identically along $\mathcal{Y}_{b}.$
This is shown precisely as in the proof of [36, Prop 3.10] - for completeness a proof is given in Step 1 in Section 6.3 below.
Step2: The inequality holds in the case of Step 1.
First observe that
(6.2) $\mathcal{M}_{\mathcal{Y}}(\mathcal{L}^{\prime})-\mathcal{M}_{\mathcal{X}}(\mathcal{L})=\mathcal{L}^{\prime}\cdot\left(\mathcal{K}_{\mathcal{Y}/\mathcal{X}}-\frac{1}{2}E\right).$
Indeed, rewriting
$\mathcal{M}_{\mathcal{X}}(\mathcal{L})=-\frac{\mathcal{L}^{2}}{2}+\cdot\mathcal{L}\cdot(\mathcal{L}+\mathcal{K}_{\mathcal{X}})$
(and likewise for $(\mathcal{Y},\mathcal{L}^{\prime})$) the left hand side in formula 6.2 may be expressed as
$\frac{p^{*}\mathcal{L}^{2}-\mathcal{L}^{\prime 2}}{2}+\mathcal{L}^{\prime}\cdot(\mathcal{K}_{\mathcal{Y}/\mathcal{X}}-E)=\frac{E\cdot(p^{*}\mathcal{L}+\mathcal{L}^{\prime})}{2}+\mathcal{L}^{\prime}\cdot(\mathcal{K}_{\mathcal{Y}/\mathcal{X}}-E).$
Since $E\cdot p^{*}\mathcal{L}=0$ this proves formula 6.2.
Since $\mathcal{L}^{\prime}$ is relatively semi-ample it will thus be enough to show that the vertical exceptional divisor $\mathcal{K}_{\mathcal{Y}/\mathcal{X}}-\frac{1}{2}E$ is effective. This means, by the assumption on $\alpha(\mathcal{X})(:=\alpha(\mathcal{X},0)),$ that it is enough to show that
(6.3) $\mathcal{K}_{\mathcal{Y}/\mathcal{X}}-\alpha(\mathcal{X})E\geq 0$
To fix ideas first assume that $\pi(\mathcal{Z})$ is supported on a single point that we denote by $b.$ By Step 1, we can express $s_{b}=p^{*}s$ for a global section $s$ of $\mathcal{L}\rightarrow\mathcal{X}$ whose zero-divisor $F$ does not vanish identically on $\mathcal{X}_{b}$ and such that $p^{*}F-E$ is effective. Since $F$ is a contender for the inf defining $\alpha(\mathcal{X})$ we have
$\alpha(\mathcal{X})\leq\text{lct}(\mathcal{X},\mathcal{X}_{b};\,F),$
Next, since $p^{*}F\geq E$ it follows from Lemma 6.1 that
$\text{lct}(\mathcal{X},\mathcal{X}_{b};\,F)\leq\inf_{i}\frac{a_{i}+1-b_{i}}{c_{i}}$
where $i$ runs over the $p-$exceptional irreducible prime divisors $E_{i}$ of $\mathcal{Y}$ and $a_{i},b_{i}$ and $c_{i}$ denote the order of vanishing along $E_{i}$ of $\mathcal{K}_{\mathcal{Y}/\mathcal{X}},$ $\mathcal{Y}_{b}$ and $E$ respectively. Note that $c_{i}>0$ (since the support of $E$ contains the support of all $p-$exceptional divisors) and $b_{i}\geq 1$ (since $\mathcal{Z}$ is assumed to be supported in $\mathcal{X}_{b}).$ Thus
$\alpha(\mathcal{X})\leq\inf_{i}\frac{a_{i}+1-b_{i}}{c_{i}}\leq\inf_{i}\frac{a_{i}}{c_{i}},$
giving
$\mathcal{K}_{\mathcal{Y}/\mathcal{X}}-\alpha(\mathcal{X})E\geq\sum_{i}a_{i}E_{i}-\left(\min_{j}\frac{a_{j}}{c_{j}}\right)E_{i}=\sum_{i}\left(\frac{a_{i}}{c_{i}}-\left(\min_{j}\frac{a_{j}}{c_{j}}\right)\right)c_{i}E_{i}\geq 0,$
which proves 6.3. Finally, consider the general case when the support of $\pi(\mathcal{Z})$ consists of a finite number of points $b_{m}$ in $\mathcal{B}.$ We then split the vertical divisors $\mathcal{K}_{\mathcal{Y}/\mathcal{X}}$ and $E$ into the components $\mathcal{K}_{\mathcal{Y}/\mathcal{X}}^{(m)}$ and $E^{(m)}$ over $b_{m}:$
$\mathcal{K}_{\mathcal{Y}/\mathcal{X}}-\alpha(\mathcal{X})E=\sum_{m}\mathcal{K}_{\mathcal{Y}/\mathcal{X}}^{(m)}-\alpha(\mathcal{X})E^{(m)}$
and apply the previous bound for each fixed $m$ (with $b$ replaced by $b_{m})$ to get that $\mathcal{K}_{\mathcal{Y}/\mathcal{X}}^{(m)}-\alpha(\mathcal{X})E^{(m)}\geq 0$ and thus $\mathcal{K}_{\mathcal{Y}/\mathcal{X}}-\alpha(\mathcal{X})E\geq 0,$ as desired.
Proof for log pairs
Just as in the previous case it is enough to consider the special case of Step 2. In this case formula 6.2 readily generalizes to
$\mathcal{M}_{(\mathcal{Y},q^{*}\mathcal{D}^{\prime})}(\mathcal{L}^{\prime})-\mathcal{M}_{(\mathcal{X},\mathcal{D})}(\mathcal{L})=\mathcal{L}^{\prime}\cdot\left(\mathcal{D}^{\prime}-p^{*}\mathcal{D}+\mathcal{K}_{\mathcal{Y}/\mathcal{X}}-\frac{1}{2}E\right).$
As before it will thus be enough to show that
(6.4) $\mathcal{K}_{\mathcal{Y}/\mathcal{X}}+q^{*}\mathcal{D}^{\prime}-p^{*}\mathcal{D}-\alpha(\mathcal{X},\mathcal{D})E\geq 0.$
To simplify the exposition we will assume that $\pi(\mathcal{Z})$ is a single closed point in $\mathcal{B},$ denoted by $b$ (the general case is shown in a similar way by decomposing wrt the components of $\pi(\mathcal{Z})$ as above). By the definition of $\alpha(\mathcal{X},\mathcal{D})$
(6.5) $\alpha(\mathcal{X},\mathcal{D})\leq\text{lct}(\mathcal{X},\mathcal{D}+\mathcal{X}_{b};\,F),$
Next, since $p^{*}F-E$ is effective, i.e. $p^{*}F\geq E$ Lemma 6.1 yields
$\text{lct}(\mathcal{X},\mathcal{D}+\mathcal{X}_{b};\,F)\leq\inf_{i}\frac{a_{i}+1-d_{i}-b_{i}}{c_{i}},$
where $a_{i},b_{i},$ $c_{i}$ and $d_{i}$ are the order of vanishing along $E_{i}$ of $\mathcal{K}_{\mathcal{Y}/\mathcal{X}},$ $\mathcal{Y}_{b}$, $E$ and $p^{*}\mathcal{D}$ respectively. In particular, $b_{i}\geq 1$ since $\mathcal{Z}$ is supported in $\mathcal{X}_{b},$ Hence,
$\alpha(\mathcal{X},\mathcal{D})\leq\inf_{i}\frac{a_{i}+1-d_{i}-b_{i}}{c_{i}}.$
Next, we may decompose
$p^{*}\mathcal{D}=\left(p^{*}\mathcal{D}\right)_{\text{hor}}+(p^{*}\mathcal{D})_{\text{ex}},$
where $\left(p^{*}\mathcal{D}\right)_{\text{hor}}$ is the horizontal divisor obtained as the proper transform of the horizontal divisor $\mathcal{D}$ and $(p^{*}\mathcal{D})_{\text{ex}}$ is $p-$exceptional. By 6.5 $\mathcal{K}_{\mathcal{Y}/\mathcal{X}}-(p^{*}\mathcal{D})_{\text{ex}}-\alpha E_{\text{ex}}\geq$
$\geq\sum_{i}(a_{i}-d_{i})E_{i}-\left(\min_{j}\frac{a_{j}-b_{j}+1-d_{j}}{c_{j}}\right)c_{i}E_{i}\geq\sum_{i}\left(\frac{(a_{i}-d_{i})}{c_{i}}E_{i}-\left(\min_{j}\frac{a_{j}-d_{j}}{c_{j}}\right)\right)c_{i}E_{i}\geq 0$
using that $b_{j}\geq 1.$ Hence,
$\mathcal{K}_{\mathcal{Y}/\mathcal{X}}+\mathcal{D}^{\prime}-p^{*}\mathcal{D}-\alpha(\mathcal{X},\mathcal{D})E_{\text{ex}}\geq\mathcal{D}^{\prime}-\left(p^{*}\mathcal{D}\right)_{\text{hor}}.$
But, since both $(\mathcal{X}^{\prime},\mathcal{D}^{\prime})$ and $(\mathcal{X},\mathcal{D})$ are models for $(X,D)$ and $\mathcal{D}$ is assumed horizontal it follows that $q^{*}\mathcal{D}^{\prime}-\left(p^{*}\mathcal{D}\right)_{\text{hor}}$ is an effective vertical divisor and hence
$q^{*}\mathcal{D}^{\prime}-\left(p^{*}\mathcal{D}\right)_{\text{hor}}\geq 0,$
which concludes the proof of the inequality 6.4.
Lemma 6.3.
Assume that $(X,D)$ is a K-semistable log Fano curve over $\mathbb{C}.$ For the canonical model $(\mathbb{P}_{\mathbb{Z}}^{1},\mathcal{D})$ of $(X,D)$
$\alpha(\mathcal{X},\mathcal{D})\geq 1/2.$
Proof.
By inversion of adjunction on surfaces over excellent schemes [46]
[…]
if $F$ does not contain the support of the divisor $\mathcal{X}_{b}.$ In the present case $\mathcal{X}_{b}=\mathbb{P}_{\mathbb{F}_{b}}^{1},$ where $b$ has been identified with a prime number and $\mathbb{F}_{b}$ denotes the field with $b$ elements. Decomposing
$D=\sum w_{i}D_{i}$
the K-semistability assumption is, by 4.2, equivalent to the condition
(6.6) $w_{j}\leq\frac{1}{2}\sum_{i}w_{i},\,\,\,\forall j.$
We recall that for any curve $C$ over a perfect field (here taken to be $\mathbb{P}_{\mathbb{F}_{b}}^{1}$) an effective $\mathbb{Q}-$divisor $F$ on $C$ is lc iff all its coefficients are less then are equal to one [27, 46]. Since $-K_{\mathbb{P}_{\mathbb{F}_{b}}^{1}}$ is linearly equivalent to $\mathcal{O}(2)$ and $D_{i|\mathcal{X}_{b}}$ is linearly equivalent to $\mathcal{O}(1)$ it thus follows from the weight condition 6.6 that $\text{lct}(\mathcal{X},\mathcal{D}_{|\mathcal{X}_{b}};F_{|\mathcal{X}_{b}})\geq 1/2.$ Indeed, it is enough to consider the case when $F=(2-\sum_{i}w_{i})[x],$ where $[x]$ denotes the prime divisor on $\mathbb{P}_{\mathbb{F}_{b}}^{1}$ corresponding to a closed point $x$ in $\mathbb{P}_{\mathbb{F}_{b}}^{1}.$ Then
$\frac{1}{2}F+\mathcal{D}_{|\mathcal{X}_{b}}=(1-\frac{1}{2}\sum_{i}w_{i})[x]+\sum w_{i}[x_{i}],$
In the case that $x\neq x_{i}$ for any $i$ the coefficients of $F/2+\mathcal{D}_{|\mathcal{X}_{b}}$ are indeed less than or equal to $1,$ since $w_{i}\in[0,1].$ Moreover, if $x=x_{j}$ then the coefficient of index $j$ equals $(1-\frac{1}{2}\sum_{i}w_{i})+w_{j}$ which is less than are equal to $1,$ by the weight condition 6.6.
We will also need the following lemma, shown precisely as in the case when $\mathcal{D}=0$ considered in [1, Prop 5.3]. ∎
Lemma 6.4.
Let $(\mathcal{X},\mathcal{D};\mathcal{L})$ be a polarized arithmetic log surfaces $(\mathcal{X},\mathcal{D};\mathcal{L})$ such that the complexification $(X,\Delta)$ of $(\mathcal{X},\mathcal{D})$ is a Fano variety and $\mathcal{L}\otimes\mathbb{C}=-K_{(X,\Delta)}.$ A metric realizes the infimum
$\inf_{\left\|\cdot\right\|}\mathcal{M}_{(\mathcal{X},\mathcal{D})}(\mathcal{L},\left\|\cdot\right\|)$
over all locally bounded metrics on $-(K_{(X,D)})$ with positive curvature current iff the metric is a log Kähler-Einstein metric. In particular, in the case when $D=0$ any minimizer coincides with the Fubini-Study metric up to the application of an automorphism of $X$ and a scaling of the metric. Moreover,
$\inf_{\left\|\cdot\right\|}\mathcal{M}_{(\mathbb{P}_{\mathbb{Z}}^{1},\mathcal{D})}\left(-\mathcal{K}_{(\mathbb{P}_{\mathbb{Z}}^{1},\mathcal{D})},\left\|\cdot\right\|\right)=-\sup_{\left\|\cdot\right\|}\frac{1}{2}\left(-\mathcal{K}_{(\mathbb{P}_{\mathbb{Z}}^{1},\mathcal{D})},\left\|\cdot\right\|\right)^{2}$
where the sup in the right hand side is restricted to volume-normalized metrics.
6.3. Conclusion of the proof of Theorem 1.2 and Corollary 1.6
Combining the previous first two lemmas immediately yields
$\mathcal{M}_{(\mathcal{X}^{\prime},\mathcal{D}^{\prime})}(\overline{\mathcal{L}^{\prime}})\geq\mathcal{M}_{(\mathbb{P}_{\mathbb{Z}}^{1},\mathcal{D})}(\overline{-\mathcal{K}_{(\mathbb{P}_{\mathbb{Z}}^{1},\mathcal{D})}}).$
Applying the third lemma above thus gives
$\mathcal{M}_{(\mathcal{X}^{\prime},\mathcal{D}^{\prime})}(\overline{\mathcal{L}^{\prime}})\geq-\sup_{\left\|\cdot\right\|}\left(\frac{1}{2}\left(-\mathcal{K}_{(\mathbb{P}_{\mathbb{Z}}^{1},\mathcal{D})},\left\|\cdot\right\|\right)^{2}\right)$
where the infimum in the left hand side is restricted to volume-normalized metrics. Invoking Theorem 1.3 and using that the Fubini-Study metric is a minimizer when $\mathcal{D}=0$ (by Lemma 6.4) thus proves the inequality in Theorem 1.5 and it corollary. Moreover, Theorem 1.3 implies that the inequality is strict, as soon as $D$ is non-trivial.
6.3.1. The equality case
Consider now the case of equality in Theorem 1.5 (and, as a consequence, $D=0):$
(6.7) $\mathcal{M}_{(\mathcal{X}^{\prime},\mathcal{D}^{\prime})}(\overline{\mathcal{L}^{\prime}})=\mathcal{M}_{(\mathbb{P}_{\mathbb{Z}}^{1},0)}(\overline{-\mathcal{K}_{\mathbb{P}_{\mathbb{Z}}^{1}}})$
where, in the rhs, $-\mathcal{K}_{\mathbb{P}_{\mathbb{Z}}^{1}}$ is endowed with the Fubini-Study metric. By the minimizing property in Lemma 6.1, when $D=0,$ the metric on $\mathcal{L}$ coincides with the Fubini-Study metric up to the application of an automorphism of $X$ and scaling of the metric. All that remains is to show is thus that $(\mathcal{X}^{\prime},\mathcal{D}^{\prime})$ is isomorphic to $(\mathbb{P}_{\mathbb{Z}}^{1},0).$ To this end first note that since $\mathcal{X}(=\mathbb{P}_{\mathbb{Z}}^{1})$ and $\mathcal{X}^{\prime}$ have the same generic fiber they are birationally equivalent. Thus, there exists a normal variety $\mathcal{Y},$ which is flat and projective over $\mathcal{B},$ dominating both $\mathcal{X}$ and $\mathcal{X}^{\prime},$ with birational morphisms
(6.8) $p:\mathcal{Y}\rightarrow\mathcal{X},\,\,\,q:\mathcal{Y}\rightarrow\mathcal{X}^{\prime}$
which are the identity over the generic point in $\mathcal{B}$ (a concrete construction is given in Step 1 below). It will thus be enough to show that the equality 6.7 implies that $p$ can be taken to be an isomorphism. Indeed, if $p$ is an isomorphism we get a birational morphism $q$ from $\mathbb{P}_{\mathbb{Z}}^{1}$ to $\mathcal{X}^{\prime}$ and any such morphism is an isomorphism (since the fibers of $\mathbb{P}_{\mathbb{Z}}^{1}$ over $\mathcal{B}$ are all reduced and irreducible). Moreover, when $\mathcal{X}^{\prime}$ is equal to $\mathbb{P}_{\mathbb{Z}}^{1}$ any $\mathcal{L}$ whose complexification equals $-K_{\mathbb{P}^{1}}$ is isomorphic to $-\mathcal{K}_{\mathbb{P}_{\mathbb{Z}}^{1}}$ and the components of any divisor $\mathcal{D}^{\prime}$ on $\mathcal{X}^{\prime}$ whose complexification is trivial are fibers $\mathcal{X}_{b_{i}}$ of $\mathbb{P}_{\mathbb{Z}}^{1}$ (using again that the fibers of $\mathbb{P}_{\mathbb{Z}}^{1}$ over $\mathcal{B}$ are all reduced and irreducible). Hence, the assumed equality 6.7 implies, since $\mathcal{X}_{b_{i}}^{2}=0$ and $-\mathcal{K}_{\mathbb{P}_{\mathbb{Z}}^{1}}$ is relatively ample that $\mathcal{D}^{\prime}$ is trivial, i.e. $\mathcal{D}^{\prime}=0.$
Thus all that remains is to show that the assumed equality in formula 6.7 implies that the morphism $p:\mathcal{Y}\rightarrow\mathcal{X}$ (in formula 6.8) can be taken to be an isomorphism.
Step 1: In the case of arithmetic surfaces $p:\,\mathcal{Y}\rightarrow\mathcal{X}$ can be taken as the successive blow-ups of $\mathcal{X}$ along a finite number of closed points $x_{i}$ in regular surfaces $\mathcal{X}_{i}$ and there exists a $p-$exceptional and $p-$ample divisor $E$ on $\mathcal{Y}$ and a morphism $q$ from $\mathcal{Y}$ to $\mathcal{X}$ such that $q^{*}\mathcal{L}^{\prime}=p^{*}\mathcal{L}-E.$ In particular, $\mathcal{M}_{\mathcal{X}^{\prime}}(\mathcal{L}^{\prime})=\mathcal{M}_{\mathcal{Y}}(p^{*}\mathcal{L}-E).$
This is shown as in the beginning of the proof of [36, Prop 3.10], as next explained. First note that since $\mathcal{X}$ and $\mathcal{X}^{\prime}$ have the same generic fiber they are birationally equivalent. Since $\mathcal{X}$ is normal this means that there exists a morphism $h:U\rightarrow\mathcal{X}^{\prime}$ from a Zariski open subset $U$ in $\mathcal{X}$ of codimension two. As a consequence, $h^{*}\mathcal{L}^{\prime}$ extends to a $\mathbb{Q}-$line bundle $\mathcal{L}^{\prime\prime}$ on $\mathcal{X}$ coinciding with $-\mathcal{K}_{\mathcal{X}}$ on the generic fiber. Since $\mathcal{X}=\mathbb{P}_{\mathbb{Z}}^{1}$ this implies that $\mathcal{L}^{\prime\prime}$ is isomorphic to $-\mathcal{K}_{\mathcal{X}}$ (using, for example, that $\pi:\mathcal{X}\rightarrow\text{Spec $\mathbb{Z}$}$ has reduced irreducible fibers). Now fix a positive integer $k$ such that $k\mathcal{L}^{\prime}$ is a relatively very ample line bundle and take a basis $s^{\prime}_{i}$ in the free $\mathbb{Z}-$module $H^{0}(\mathcal{X}^{\prime},k\mathcal{L}^{\prime}$). Then $s_{i}:=h^{*}s_{i}$ extends, since $\mathcal{X}$ is normal, to a unique element in $H^{0}(\mathcal{X},k\mathcal{L}).$ Denote by $\mathcal{J}$ the ideal sheaf on $\mathcal{X}$ generated by the sections $s_{i}.$ Since $\mathcal{X}$ is a regular surface we get after successive blow-ups (as stated in Step 1) a morphism $p:\mathcal{Y}\rightarrow\mathcal{X}$ from a regular surface $\mathcal{Y}$ to $\mathcal{X}$ with the property that $p^{*}\mathcal{J}$ defines an effective $p-$exceptional divisor $E_{k}$ on $\mathcal{Y}$ (using that $\mathbb{Z}$ is an excellent ring) [53]. Set
$E:=k^{-1}E_{k}\,\,\,\,(E_{k}:=p^{*}\mathcal{J})$
By construction, $E_{k}$ is $p-$ample,
(6.9) $H^{0}(\mathcal{Y},kp^{*}\mathcal{L}-E_{k})\cong H^{0}(\mathcal{X},k\mathcal{L}\otimes\mathcal{J})\cong H^{0}(\mathcal{X}^{\prime},k\mathcal{L}^{\prime})$
and the global sections of $kp^{*}\mathcal{L}-E_{k}$ induce a morphism $q$ to $\mathcal{X}^{\prime}$ such that $q^{*}\mathcal{L}^{\prime}=p^{*}\mathcal{L}-E.$ Finally, note that
$\mathcal{M}_{\mathcal{X}^{\prime}}(\mathcal{L}^{\prime})=\mathcal{M}_{\mathcal{Y}}(q^{*}\mathcal{L}^{\prime}),$
as follows directly from the fact that $p$ is an isomorphism between Zariski open subsets of $\mathcal{X}^{\prime}$ and $\mathcal{Y}$ and, as a consequence, the $\mathbb{Q}-$line bundle $q^{*}\mathcal{L}^{\prime}$ is trivial on the support of the divisor $q^{*}\mathcal{K}_{\mathcal{X}}-\mathcal{K}_{\mathcal{Y}}.$
Step 2: $\mathcal{M}_{\mathcal{Y}}(p^{*}\mathcal{L}-E)=\mathcal{M}_{\mathcal{X}}(\mathcal{L})\implies p\text{ is an isomorphism, when $\mathcal{X}=\mathbb{P}_{\mathbb{Z}}^{1}$}$
Replacing $q^{*}\mathcal{L}^{\prime}$ with $p^{*}\mathcal{L}-E$ in formula 6.2 yields, since $\mathcal{M}_{\mathcal{Y}}(p^{*}\mathcal{L}-E)=\mathcal{M}_{\mathcal{X}}(\mathcal{L}),$
$(p^{*}\mathcal{L}-E)\cdot\left(\mathcal{K}_{\mathcal{Y}/\mathcal{X}}-\frac{1}{2}E\right)=0.$
It follows, since, by construction, $p^{*}\mathcal{L}-E$ is $p-$ample, that
(6.10) $\mathcal{K}_{\mathcal{Y}/\mathcal{X}}=\frac{1}{2}E.$
Now, since $p:\,\mathcal{Y}\rightarrow\mathcal{X}$ is the blow-up along a finite number of closed points $x_{i}$ in regular surfaces $\mathcal{X}_{i},$
(6.11) $\mathcal{K}_{\mathcal{Y}/\mathcal{X}}=\sum c_{i}E_{i},\,\,\,c_{i}\geq 1$
where the sum runs over all prime $p-$exceptional divisors $E_{i}.$ Hence,
$E=\sum_{i}2c_{i}E_{i}\geq\sum_{i}2E_{i}.$
But this contradicts the isomorphisms 6.9, if the number of points $x_{i}$ is non-zero. Indeed, denote by $E_{1}$ the strict transform of the $p-$exceptional divisor on $\mathcal{Y}$ induced from the exceptional divisor on the first point $x_{1}$ blown-up on $\mathcal{X}(=\mathbb{P}_{\mathbb{Z}}^{1}).$ Then it it follows from the previous inequality and the construction of $E$ that the restriction of the ideal sheaf $\mathcal{J}$ on $\mathcal{X}$ to a neighbourhood of $x_{1}$ in the fiber $\mathcal{X}_{\pi(x_{1})}$ is contained in the $2k$th power $\mathfrak{m}_{x_{1}}^{2k}$ of the maximal ideal $\mathfrak{m}_{x_{1}}$ on $\mathcal{X}_{\pi(x_{1})}$ defined by the point $x_{1}.$ But, in general, for $\mathcal{X}=\mathbb{P}_{\mathbb{Z}}^{1},$ the line bundle $k\mathcal{L}_{|\mathcal{X}_{\pi(x)}}\otimes\mathfrak{m}_{x}^{2k}$ on $\mathcal{X}_{\pi(x)}$ is trivial for any closed point $x$ on $\mathcal{X}$ (since $\mathcal{L}_{|\mathcal{X}_{b}}:=-K_{\mathcal{X}_{b}}=\mathcal{O}_{\mathbb{P}_{\mathbb{F}_{b}}^{1}}(2)).$ But this contradicts the isomorphism 6.9, since $\mathcal{L}^{\prime}$ is relatively ample. Hence, the number of points $x_{i}$ must be zero, as desired.
Combing these two steps thus concludes, as discussed above, the proof of Theorem 1.5. Finally, Corollary 1.6 can be deduced from Theorem 1.5 using a generalization of Lemma 6.4. But here we instead proceeds as follows. Given an arithmetic log Fano surface $(\mathcal{X},\mathcal{D})$ set $\mathcal{L}:=-\mathcal{K}_{\mathcal{X}}$ and endow $\mathcal{L}$ and $-\mathcal{K}_{\mathcal{X}}$ with the same metric induced from a volume-normalized metric on $-K_{X}$ with positive curvature current. Then, by definition 6.1,
$-\frac{1}{2}\overline{\mathcal{K}}_{(\mathcal{X},\mathcal{D})}^{2}=\mathcal{M}_{(\mathcal{X},\mathcal{D})}(\mathcal{L}).$
Hence, combining Step one and Step two above yields
$-\frac{1}{2}\overline{\mathcal{K}}_{(\mathcal{X},\mathcal{D})}^{2}\geq\mathcal{M}_{(\mathbb{P}_{\mathbb{Z}}^{1},0)}(\overline{-\mathcal{K}_{\mathbb{P}_{\mathbb{Z}}^{1}}})=-\frac{1}{2}(\overline{-\mathcal{K}_{\mathbb{P}_{\mathbb{Z}}^{1}}})^{2}$
and the equality case is deduced precisely as before.
7. Appendix
In the proof of Lemma 5.3 we used the following result (applied to $X=\mathbb{P}_{\mathbb{C}}^{n}$).
Lemma 7.1.
Let $X$ be a Fano manifold and $V$ a holomorphic vector field on $X.$ Denote by $F_{t}$ the flow of the real part of $V$ on $X$ at time $t$ and by $(F_{t}^{V})^{*}\phi$ its action on a given continuous metric $\phi$ on $-K_{X}$ with positive curvature current. If $X$ admits a Kähler-Einstein metric, then
$\frac{d}{dt}\mathcal{E}(F_{t}^{*}\phi,\psi_{0})=0$
for any fixed metric $\psi_{0}$ on $-K_{X}.$
Proof.
This is well-known and essentially goes back to [24], but for the convenience of the reader we provide a proof in the spirit of the present paper and its precursor [2]. Consider the Ding functional on the space of all continuous metrics on $-K_{X}$ with positive curvature current, defined by
$\mathcal{D}(\phi):=-\frac{n!}{(-K_{X})^{n}}\mathcal{E}(\phi)-\log\int_{X}\mu_{\phi},$
where $\mu_{\phi}$ is the measure on $X$ corresponding to the metric $\phi$ (see Section 2.1, with $\Delta=0)$ and $\mathcal{E}(\phi)$ is a shorthand for the functional $\mathcal{E}(\phi,\psi_{0})$ defined in formula 2.7. Since, $\mu_{F^{*}\phi}=F^{*}\mu_{\phi}$ for any biholomorphism $F$ of $X$ it follows that $\mathcal{E}(F_{t}^{*}\phi)$ and $\mathcal{D}(F_{t}{}^{*}\phi)$ have the same derivative. Moreover, in general, the function $t\mapsto\mathcal{D}(F_{t}^{*}\phi)$ is linear. Indeed, its derivative is the Futaki invariant of $V$ (see the claim on page 73 in [48], where $\mathcal{D}$ is denoted by $F_{\omega}$ ). Hence, all that remains is to verify that $\mathcal{D}$ is bounded from below (since then $t\mapsto\mathcal{D}(F_{t}^{*}\phi)$ must be constant). But this follows from the existence of a Kähler-Einstein metric, since such a metric minimizes $\mathcal{D},$ as recalled in [2, Section 2.3]. ∎